### v1

# PROTOCOL FOR AN OBSERVATIONAL STUDY ON THE EFFECTS OF COMBINATIONS OF ADVERSE CHILDHOOD EXPERIENCES ON ADULT DEPRESSION

Abstract.
Adverse childhood experiences (ACEs) have been linked to a wide range of negative health outcomes in adulthood. However, few studies have investigated what specific combinations of ACEs most substantially impact mental health. In this article, we provide the protocol for our observational study of the effects of combinations of ACEs on adult depression. We use data from the 2023 Behavioral Risk Factor Surveillance System (BRFSS) to assess these effects. We will evaluate the replicability of our findings by splitting the sample into two discrete subpopulations of individuals. We employ data turnover for this analysis, enabling a single team of statisticians and domain experts to collaboratively evaluate the strength of evidence, and also integrating both qualitative and quantitative insights from exploratory data analysis. We outline our analysis plan using this method and conclude with a brief discussion of several specifics for our study.
Corresponding authors:
William Bekerman, bekerman@wharton.upenn.edu; Ruizhe Zhang, zhangrz22@m.fudan.edu.cn
1. Introduction
Adverse childhood experiences (ACEs) describe specific categories of early life hardship or harm that occur before an individual reaches adulthood. Although these can take various forms, they are typically categorized into ten types involving abuse (physical, emotional, or sexual) and neglect (physical or emotional), as well as household challenges including mental illness, substance abuse, intimate partner violence, parental absence due to divorce or separation, or incarceration. Initially investigated in a landmark study conducted by the Centers for Disease Control and Prevention (CDC) and Kaiser Permanente, ACEs are highly prevalent and exhibit a strong dose-response relationship with numerous negative health and social outcomes in adulthood (Felitti et al. [1998]; Dube et al. [2003]; Anda et al. [2006]). Exposure to individual ACEs has been repeatedly associated with a range of adverse physical and emotional health consequences later in life. One study indicated that childhood emotional abuse increases the risk of lifetime depressive disorders, with adjusted odds ratios (ORs) of 2.7 for women and 2.5 for men (Chapman et al. [2004]). Depression also appears to be closely related to having experienced parental alcohol abuse (Anda et al. [2002]). Moreover, physical and verbal abuse have been found to be strongly associated with body weight issues and obesity (Williamson et al. [2002]).
Unfortunately, when children are exposed to one type of adversity, there is a high likelihood of exposure to additional ACEs (Dong et al. [2004]; Felitti [2002]). Numerous large-scale investigations have confirmed that these detrimental experiences can have a cumulative effect: each additional type of adversity elevates the probability of unfavorable outcomes (Felitti et al. [1998]; Chapman et al. [2004]). Accordingly, previous literature has examined closely the number of categories of ACEs, often referred to as ACE score. For instance, one study using data from 14 states revealed substantial variation across states in the percentage change in health outcomes for those exposed to at least four ACEs compared to those with fewer than four ACEs (Waehrer et al. [2020]). Physically, individuals with four or more ACEs face elevated risks of heart disease, liver disease, kidney disease, asthma, Alzheimer’s disease, stroke, metabolic disorders, obesity, diabetes, and certain cancers (for Disease Control and Prevention [2017]; for Youth Wellness [2014]; Merrick [2019]; Hughes et al. [2017]; Waehrer et al. [2020]). On the emotional and psychological side, ACE score strongly correlates with mental health issues such as depression, anxiety, and suicidal ideation (Hillis et al. [2004]; Chapman et al. [2004]). Meta-analyses further reinforce that an accumulation of at least 4 ACEs correlates with higher risks of chronic disease, poor mental health, and negative social outcomes, demonstrating the need for strategies to prevent ACEs and lessen their long-term repercussions (Hughes et al. [2017]).
A recent study challenged the overall ACE score and advised to break down the measure back into specific items (Fitzgerald and Bishop [2024]). Gaps also remain in our understanding of how the accumulation of ACEs influences health. We are therefore interested in which categories of ACEs—and which specific combinations of ACEs (e.g., physical abuse + sexual abuse, or physical abuse + emotional neglect + substance abuse)—are most harmful and warrant primary care. We focus our analysis on depression, one of the most prevalent and burdensome mental health conditions globally. Consequently, our focus is to evaluate the effect of different combinations of specific ACE items on depression and identifying a higher-risk group of sets of ACEs.
We will examine the impacts of combinations of ACEs on adult depression using data from the Behavioral Risk Factor Surveillance System (BRFSS) 2023, an ongoing health-related telephone survey system administered by the CDC with a random sample of over 400,000 individuals across the United States in 2023. This data includes detailed information from ten states that administered the CDC’s Adverse Childhood Experiences module in 2023. We can obtain all ten ACE categories proposed in the CDC-Kaiser study (Felitti et al. [1998]), and there are over 50,000 individuals with complete records of all ACE items. Among them, around 17,000 reported zero ACEs while the rest were exposed to at least one ACE. This large sample size will help facilitate powerful data analysis and explore the influencing mechanism of different ACEs combinations. The BRFSS also collected measurements of variables that may confound the relationship between exposure to ACEs and depression, such as those related to personal demographics. We will utilize those variables to form separate strata for treated and control individuals, enabling us to compare the outcomes of similar individuals in hopes of ameliorating confounding biases.
We employ a novel statistical design called data turnover (Bekerman et al. in preparation, see Bekerman et al. [2024a] for a description) to conduct this analysis. By integrating both qualitative and quantitative insights from data exploration, data turnover enables a single team of statisticians and domain experts to evaluate evidence from multiple data splits. Exploratory data analysis (EDA) allows us to inspect the data before making potentially questionable assumptions, detect anomalies or outliers that do not represent natural variation in the data, and formulate new hypotheses. It also helps identify variables that may not measure what we initially intended, indicating the need for a better measure or refined, data-informed hypotheses (Bekerman et al. [2024b]). Beyond exploration, replication is critical in observational studies for reinforcing the credibility of findings and sustaining trust in scientific knowledge. Previous work has combined some—but not all—aspects of exploration and replication within the same observational study. For instance, some methods allow for exploration without replication (e.g., Cox [1975]; Heller et al. [2009]), whereas others enable replication without exploration (e.g., Zhao et al. [2018]; Karmakar et al. [2019]). There is also research that supports both exploration and replication but requires two independent investigative teams (Roy et al. [2022]). In contrast, data turnover makes it possible to conduct both exploration and replication using a single research group.
In observational studies like ours, where treatment assignment is not under our direct control, obtaining consistent findings across multiple groups that may have different assignment mechanisms reinforces the evidence that the treatment itself is truly responsible for the observed effects (Rosenbaum [2001]; Rosenbaum [2015]). By analyzing data from two distinct subpopulations of BRFSS 2023 participants—residing in states that voted for the Democratic candidate in the 2024 US presidential election (”blue states”) and those residing in states that voted for the Republican candidate (”red states”) —we can assess the replicability of our findings. We split our data into blue states and red states, recognizing that these political orientations often shape distinct social, economic, and health care policies. A study focusing on population in East and West Germany has shown that socio-political context is a determinant of childhood maltreatment (Ulke et al. [2021]). In our context, for example, blue states may allocate more resources toward public health programs (e.g., adopt medicaid expansion) or have broader social safety nets, whereas red states may emphasize different strategies for healthcare funding and child protective services. Such policy and cultural variations can influence both the prevalence of ACEs and the mechanisms by which individuals are exposed to them, potentially resulting in different patterns of ACEs across these two groups. Demonstrating consistent results in both blue and red states thus enhances confidence that our findings are robust across varied socio-political contexts. By employing data turnover, we can predetermine how one subgroup will guide analysis of the other, then apply exploratory data analysis to the second subgroup to inform how we draw inferences about the first. Any outcome that shows a significant effect in both groups is defined as a “replicable” finding. (Bogomolov and Heller [2023]). In Section 5, we detail how we apply this non-random cross-screening method to detect replicable outcomes and test the global null hypothesis, i.e., identifying outcomes that are affected in at least one of the two subpopulations.
The remainder of the document is organized as follows. We introduce the BRFSS dataset and describe our data processing in Section 2. In Section 3, we report our candidate sets of combinations. In Section 4, we outline our stratification methodology. Finally, we conclude with a description of our data analysis procedure and explanation of several specifics of implementation in Section 5.
2. The Behavioral Risk Factor Surveillance System (BRFSS)
This study utilizes data from BRFSS, an ongoing health-related telephone survey system administered by the CDC. BRFSS collects comprehensive data on a wide range of health-related behaviors, chronic health conditions, healthcare access, and preventive service usage from the non-institutionalized adult population (18 years and older). The survey employs a dual-frame sampling methodology, using both landline and cellular telephone surveys to ensure representativeness of the data. The dataset collected in 2023 is the latest data available, has 433,323 observations from 48 states, the District of Columbia, Guam, Puerto Rico, and the US Virgin Islands, and includes topics such as health status, chronic conditions, health behaviors, healthcare access, and preventive services, as well as optional modules including social determinants of health, cancer screenings, and cognitive decline.
For the purpose of our study, we focus on the 10-state [^1] Delaware, Florida, Georgia, Missouri, Nevada, New Jersey, Oregon, Rhode Island, Tennessee, Virginia. sample that administered the CDC’s ACEs module in 2023. Following previous studies utilizing the BRFSS data (Bhan et al. [2014]; Gilbert et al. [2015]; Waehrer et al. [2020]), we collapse different questions into particular indicators. More details are provided in table 1. In total, we create 10 ACEs items from the BRFSS data (see table 2), which is consistent with the categorization in CDC-Kaiser study (Felitti et al. [1998]). It is worth noting that we first follow those criteria for data pre-processing, but these steps could be adjusted when we carry out exploratory data analysis in the data turnover..
Indicator Questions Collapsed Encoding Method
Any Sexual Abuse Number of times forced to touch, be touched, or have sex with anyone at least five years older or an adult Collapsed into a single indicator for any sexual abuse
Household Substance Use Alcohol and illegal drug use in the household Combined into a single indicator for household substance use
Physical Abuse Frequency of physical abuse Encoded as 1 if the answer is “once” or “more than once”
Verbal Abuse Frequency of verbal abuse Encoded as 1 if the answer is “once” or “more than once”
Parental Divorce Response to parental divorce question Affirmative response encoded as 1
Living with a Depressed or Mentally Ill Person Response to question about living with a mentally ill person Affirmative response encoded as 1
Living with an Incarcerated Person Response to question about living with an incarcerated person Affirmative response encoded as 1
Physical Violence Between Parents Frequency of physical violence between parents Affirmative response encoded as 1
Safety Did an adult make you feel safe and protected ‘Never’ and ‘A little of the time’ encoded as 0, ‘Some of the time’, ‘Most of the time’, ‘All of the time’ encoded as 1
Basic Needs Met Did an adult make sure basic needs were met ‘Never’ and ‘A little of the time’ encoded as 0, ‘Some of the time’, ‘Most of the time’, ‘All of the time’ encoded as 1
Table 1. Data Processing Steps and Collapsed Indicators
ACE Item Variable
Live With Anyone Depressed, Mentally Ill, Or Suicidal ACEDEPRS
Substance Abuse in the Home ACESUB
Live With Anyone Who Served Time in Prison or Jail ACEPRIS
Were Your Parents Divorced/Separated ACEDIVRC
How Often Did Your Parents Beat Each Other Up ACEPUNCH
How Often Did A Parent Physically Hurt You In Any Way ACEHURT1
How Often Did A Parent Swear At You ACESWEAR
Sexual Abuse ACESEX
Did An Adult Make You Feel Safe And Protected ACEADSAF
Did An Adult Make Sure Basic Needs Were Met ACEADNED
Table 2. ACE Items in BRFSS 2023
After processing the data, we restrict our sample to those who do not have any missing values in any ten ACEs variables. We also omit those individuals who do not have the appropriate measurement for depression. For any missing family background, we directly use the imputation value provided by the BRFSS 2023. In total, there are 49,811 remaining individuals, among whom 23,506 individuals are from blue states [^2] Delaware, New Jersey, Oregon, Rhode Island, Virginia (as established in the 2024 United States presidential election) and 26,305 individuals from red states [^3] Florida, Georgia, Missouri, Nevada, Tennessee. In the part of data with individuals from blue states, 8,076 individuals do not have any ACEs while 15,512 were exposed to at least one ACE. In the part of data with individuals from blue states, 9,163 individuals do not have any ACEs while 17,231 were exposed to at least one ACE.
3. Pre-specified Candidate Sets of ACEs
We collect ten variables of ACEs from the BRFSS 2023 dataset. To examine the diverse effects of combinations of ACEs on depression, we first construct candidate sets of ACEs that we are interested in. Since previous literature has identified four ACEs as a critical threshold, beyond which the effects appear to intensify significantly, and that individuals with at least four ACEs are at an increased risk for chronic disease, poor mental health, and negative social outcomes (Hughes et al. [2017]; Waehrer et al. [2020]), we consider all quadruples of ACEs categories, all triples of ACEs categories, all pairs of ACEs categories, and all individual ACEs categories. We thus have 385 possible combinations of ACEs formed from ten candidate ACEs categories (see Appendix A).
4. Stratification
A previous study (Waehrer et al. [2020]) used BRFSS 2009 & 2012 to examine whether the relationship between ACEs and health outcomes, such as smoking, heavy drinking, and obesity, was similar across states, and controlled for baseline covariates including race/ethnicity, sex, and age. They also controlled for education, marital status and employment, which are post-treatment variables and likely to be affected by ACEs. Cox [1958] refers to these variables as “concomitant observations” and writes: “The essential point in our assumptions about these observations is that the value for any unit must be unaffected by the particular assignment of treatments to units actually used. In practice this means that either (a) the concomitant observations are taken before the assignment of treatments to units is made; or (b) the concomitant observations are made after the assignment of treatments, but before the effect of treatments has had time to develop…; or (c) we can assume from our knowledge of the nature of the concomitant observations concerned, that they are unaffected by the treatment.” If adjustments are made for variables that are affected by the treatment, then adjustments may remove part or all of the effect of the treatment, and hence estimators that adjust for a concomitant variable that has been affected by the treatment are generally biased (Rosenbaum [1984]). Therefore, we do not adjust for education, employment and marital status in our study.
Given the relatively small number of covariates, we seek to mitigate potential confounding bias by stratifying participants on their recorded baseline characteristics (race/ethnicity, age, sex) and then comparing those exposed to specific ACE combinations (the “treated” group) with those reporting no ACE exposure (the “control” group) within each stratum. Each hypothesis contrasts a particular ACE combination against having no ACEs. Individuals who have any additional ACEs beyond the combination of interest remain in the treated group. For example, if our focus is on the effect of substance abuse, participants exposed to both substance abuse and physical abuse are still classified under “substance abuse.” We conduct these stratifications separately for individuals residing in blue states and those in red states, resulting in $385\times 2$ distinct stratified analyses.
Specifically, we stratify participants by race/ethnicity, including White (non-Hispanic), Black (non-Hispanic), Hispanic, and Other (non-Hispanic); by age, with groups spanning 18-24, 25-34, 35-44, 45-54, 55-64, and 65+; and by sex, distinguishing between male and female participants. This results in a total of 48 distinct strata for both the blue and red subgroups. In the blue subgroup, the stratum with the lowest sample size contains 61 participants, with a median of 171. In the red subgroup, the stratum with the lowest sample size has 60 participants, with a median of 161. Detailed sample sizes for each stratum of both subgroups are provided in the appendix B.
5. Design of Data Analysis
Our inference will be based on stratified sets of participants in the BRFSS 2023, as described previously in Section 4. We aim to explore the effect of different combinations of ACEs in our aforementioned candidate sets on depression and also identify those with higher risk. We will deem those combinations of ACEs for which the effect of the treatment is found to be significantly increase the odds of depression in both subgroups - those residing in blue states and red states - as replicable findings. Also, we are interested in those combinations more than doubling the odds ratio of depression in only one subgroup.
5.1. Data Turnover Framework
We employ data turnover (Bekerman et al. in preparation, Bekerman et al. [2024a]) to analyze the data, ensuring control of the family-wise error rate (FWER; the chance of falsely rejecting any true null hypotheses at level $\alpha=0.05$) and facilitating flexible inference compared to automated cross-screening (Zhao et al. [2018]), by integrating both qualitative and quantitative insights from data exploration and formulating data-informed new hypotheses (e.g., Bekerman et al. [2024b]).
A brief description is as follows. We split our data into two subgroups: those individuals in our study residing in blue states and those residing in red states. The first subgroup, without data exploration, suggests a limited number of combinations of ACEs to be validated in the second subgroup based on some pre-determined selection procedure. Then, we do ad hoc data exploration on the second subgroup to choose how to conduct one-sided hypothesis tests on the first subgroup, possibly allowing for novel hypotheses to be identified and tested. In other words, both subgroups are used to screen and both subgroups are used to validate, but only one subgroup can generate novel hypotheses to be validated in the other subgroup while still maintaining the validity of our entire procedure. Because a limited number of combinations seek validation on the second subgroup, compared with directly controlling the FWER of 385 hypotheses on the whole data (eg. by using Bonferroni correction or Holm procedure), only a small fraction of the 385 potential hypotheses are considered in the validation sample.
5.2. Analysis Plan
We are interested in testing whether a combination of ACEs would increase the odds of depression compared with the reference population.
We plan to use the Cochran–Mantel–Haenszel test (CMH) in each subgroup to screen ACEs combinations. In the screening stage, we also need to define a cutoff of p-value (eg. 0.05). In a paper defining multimorbidity (Silber et al. [2018]), they chose form a 99.9% confidence interval in checking if the interval for the odds ratio was completely above 2. This is a stringent threshold to control for the number of hypotheses falsely passing the screening step. Yet, choosing this cutoff is rather heuristic given the sample sizes and number of hypotheses. A basic intuition would be: if the cutoff is too large, then almost all of the hypotheses would pass the screening stage, which would make this step meaningless; if the cutoff is too small, though only a few number of true nulls would falsely pass the screening stage, some true non-nulls would be screened out if the signals aren’t that strong.
Thus, following the structure of the BRFSS 2023, we set the number of participants in each subgroup at 1000, and set the number of hypotheses at 385, and try several cutoffs of p-value (0.001, 0.01,0.025, 0.05, 0.1), also varying the number of replicable hypotheses (1,50,100,150,200,250,300,350,385) and the signal-to-noise ratio (a sequence from 0.5 to 5, with a common difference 0.1). The result (see Figure 1) shows that whatever the number of replicable hypotheses and signal-to-noise ratio, the power using 0.001 as the cutoff is always the lowest. But there’s only trivial difference of power between choosing 0.01,0.025, 0.05 or 0.1, and none of them consistently perform the best under different values of the number of replicable hypotheses and the signal-to-noise ratio.
Figure 1. Simulation results of power with p-value cutoffs, number of replicable hypotheses and different signal-to-noise ratios.
Note: the $x$ axis represents the signal-to-noise ratio, the the $y$ axis represents the power. The facet titles in the gray bars are numbers of replicable hypotheses.
We intend to control the FWER at level $\alpha=0.05$. Therefore, we will apply Holm’s procedure (Holm [1979]) on each lists of p-values of selected hypotheses at level $\alpha/2=0.025$. Accordingly, we choose 0.025 as our cutoff in the screening stage. In other words, if the p-value of one hypothesis given by the CMH test is smaller than 0.025, then this hypothesis will pass the screening step, otherwise, this will be screened out. After selecting a list of hypotheses from subgroup 1 from the candidate sets collection, we will test them on subgroup 2. Specifically, for each selected hypothesis, we calculate the p-value using the data from subgroup 2, in the same way as we did in the screening stage on subgroup 1, and utilize Holm’s method at 0.025 to validate. Then, we will create a list of hypotheses from subgroup 2, including those selected from the candidate sets collection and generated by EDA, and validate on subgroup 1.
Moreover, we need to determine whether it is more beneficial to conduct EDA on the smaller or larger sample. Simulation results in Bekerman’s paper (in preparation) suggest that performing EDA on the smaller subgroup—in our case, participants living in blue states—is advantageous. We anticipate that incorporating data exploration within this subgroup will strengthen our study’s overall design by enabling data-informed decision-making and facilitating the emergence of new hypotheses. This flexibility may allow us to identify more complex hypotheses or yield unexpected benefits. For example, we may explore alternative ways of grouping ACE categories, refine our criteria for encoding ACE items, or investigate potential effect modification.
5.3. The high-risk group
After employing data turnover, we define the group of combinations for which the global null is rejected, that is, at least in one subgroup, as the high-risk group; in other words, the relevant combinations of ACEs will be identified as placing individuals at higher risk of depression. The group is formed as the union of validated combinations from subgroup 1 and validated combinations from subgroup 2, and simultaneously, we can tell what combinations are replicable and are only validated on one subgroup. If any pre-specified hypotheses passes four tests - passes the screening stage in the first subgroup and the validation in the second subgroup, and also passes the screening stage in the second subgroup and the validation in the first subgroup - they will be deemed as replicable. For new hypotheses formulated in the EDA process, if they are validated in the other subgroup, related combinations will be in the high-risk group.
We also provide a parsimonious list without duplicating any set whose components are already on the list as a smaller set. In other words, if a combination A in the collection is the subset of another combination B, we will remove B from the collection. For example, if both hypotheses regarding substance abuse and substance abuse + physical abuse are rejected, only substance abuse will be included in the parsimonious list. Previous literature has confirmed that ACEs can have a cumulative effect: each additional type of adversity elevates the probability of unfavorable outcomes (Felitti et al. [1998]; Chapman et al. [2004]). Therefore, given this monotonicity assumption, the validation of combination A implies that combination B increases the odds ratio of depression.
In terms of replicable findings, other than combinations simultaneously rejected in both subgroups, we also include the combination of non-duplicate components from two different combinations respectively rejected in each subgroup, if this combination has not been contained, based on the monotonicity assumption. For example, if substance abuse is rejected on subgroup 1, but substance abuse + physical abuse is not (which does not make sense in real-world scenario but may happen due to some statistical reasons), and only substance abuse + physical abuse is rejected on subgroup 2, then substance abuse + physical abuse will be deemed replicable.
The following tree diagrams (Figure 2, 3) makes things clear visually. We organize the ACE combinations in a tree structure, where each layer represents combinations of 1, 2, 3, or 4 ACE items, with each node corresponding to a unique combination. If a combination is rejected in the first subgroup, a yellow checkmark is placed at the top-left corner of the node, while a green checkmark appears at the bottom-right corner if the combination is rejected in the second subgroup. According to the monotonicity assumption, if a node is marked with either a yellow or green check, the entire subtree beneath that node inherits the same mark. Additionally, it’s important to note that identical combinations may appear in different subtrees within the same layer. If a combination node is marked in one subtree, all nodes with the same combination in other subtrees will be similarly marked. Ultimately, a combination node that is marked with checkmarks from both subgroups represents a replicable finding. In constructing the parsimonious list, we prioritize nodes from top to bottom, truncating any subtrees of replicable nodes as they are considered finalized, thus preventing redundant checks for combinations that have already been deemed replicable.
In Example 1 (Figure 2), only substance abuse is rejected in subgroup 1, and this node is marked with a yellow checkmark. In contrast, only the combination substance abuse + physical abuse is rejected in subgroup 2, with a green checkmark placed on this node. Moving to the second layer, all nodes in the subtree of substance abuse inherit the yellow checkmark. As a result, the combination substance abuse + physical abuse is marked with both checkmarks and is considered replicable. Simultaneously, the subtrees of substance abuse + physical abuse are truncated, meaning that any combinations containing substance abuse and physical abuse are no longer considered for further evaluation. In Example 2 (Figure 3), only sexual abuse is rejected in subgroup 1, while only physical abuse is rejected in subgroup 2. Following the rules of placing checkmarks, sexual abuse + verbal abuse has two checkmarks and is deemed replicable. Also, the subtrees of sexual abuse + verbal abuse are truncated.
Figure 2. Tree Diagram: Example 1
Figure 3. Tree Diagram: Example 2
Appendix A Candidate Sets of ACEs (variable names used)
We consider all quadruples of ACEs categories, all triples of ACEs categories, all pairs of ACEs categories, and all individual ACEs category and form this candidate sets of ACEs.
Table 3. Candidate sets of ACEs
Item.1 Item.2 Item.3 Item.4
1 ACESUB
2 ACEPRISN
3 ACEDIVRC
4 ACEPUNCH
5 ACEHURT1
6 ACESWEAR
7 ACESEX
8 ACEADSAF
9 ACEADNED
10 ACEDEPRS
11 ACESUB ACEPRISN
12 ACESUB ACEDIVRC
13 ACESUB ACEPUNCH
14 ACESUB ACEHURT1
15 ACESUB ACESWEAR
16 ACESUB ACESEX
17 ACESUB ACEADSAF
18 ACESUB ACEADNED
19 ACESUB ACEDEPRS
20 ACEPRISN ACEDIVRC
21 ACEPRISN ACEPUNCH
22 ACEPRISN ACEHURT1
23 ACEPRISN ACESWEAR
24 ACEPRISN ACESEX
25 ACEPRISN ACEADSAF
26 ACEPRISN ACEADNED
27 ACEPRISN ACEDEPRS
28 ACEDIVRC ACEPUNCH
29 ACEDIVRC ACEHURT1
30 ACEDIVRC ACESWEAR
31 ACEDIVRC ACESEX
32 ACEDIVRC ACEADSAF
33 ACEDIVRC ACEADNED
34 ACEDIVRC ACEDEPRS
35 ACEPUNCH ACEHURT1
36 ACEPUNCH ACESWEAR
37 ACEPUNCH ACESEX
38 ACEPUNCH ACEADSAF
39 ACEPUNCH ACEADNED
40 ACEPUNCH ACEDEPRS
41 ACEHURT1 ACESWEAR
42 ACEHURT1 ACESEX
43 ACEHURT1 ACEADSAF
44 ACEHURT1 ACEADNED
45 ACEHURT1 ACEDEPRS
46 ACESWEAR ACESEX
47 ACESWEAR ACEADSAF
48 ACESWEAR ACEADNED
49 ACESWEAR ACEDEPRS
50 ACESEX ACEADSAF
51 ACESEX ACEADNED
52 ACESEX ACEDEPRS
53 ACEADSAF ACEADNED
54 ACEADSAF ACEDEPRS
55 ACEADNED ACEDEPRS
56 ACESUB ACEPRISN ACEDIVRC
57 ACESUB ACEPRISN ACEPUNCH
58 ACESUB ACEPRISN ACEHURT1
59 ACESUB ACEPRISN ACESWEAR
60 ACESUB ACEPRISN ACESEX
61 ACESUB ACEPRISN ACEADSAF
62 ACESUB ACEPRISN ACEADNED
63 ACESUB ACEPRISN ACEDEPRS
64 ACESUB ACEDIVRC ACEPUNCH
65 ACESUB ACEDIVRC ACEHURT1
66 ACESUB ACEDIVRC ACESWEAR
67 ACESUB ACEDIVRC ACESEX
68 ACESUB ACEDIVRC ACEADSAF
69 ACESUB ACEDIVRC ACEADNED
70 ACESUB ACEDIVRC ACEDEPRS
71 ACESUB ACEPUNCH ACEHURT1
72 ACESUB ACEPUNCH ACESWEAR
73 ACESUB ACEPUNCH ACESEX
74 ACESUB ACEPUNCH ACEADSAF
75 ACESUB ACEPUNCH ACEADNED
76 ACESUB ACEPUNCH ACEDEPRS
77 ACESUB ACEHURT1 ACESWEAR
78 ACESUB ACEHURT1 ACESEX
79 ACESUB ACEHURT1 ACEADSAF
80 ACESUB ACEHURT1 ACEADNED
81 ACESUB ACEHURT1 ACEDEPRS
82 ACESUB ACESWEAR ACESEX
83 ACESUB ACESWEAR ACEADSAF
84 ACESUB ACESWEAR ACEADNED
85 ACESUB ACESWEAR ACEDEPRS
86 ACESUB ACESEX ACEADSAF
87 ACESUB ACESEX ACEADNED
88 ACESUB ACESEX ACEDEPRS
89 ACESUB ACEADSAF ACEADNED
90 ACESUB ACEADSAF ACEDEPRS
91 ACESUB ACEADNED ACEDEPRS
92 ACEPRISN ACEDIVRC ACEPUNCH
93 ACEPRISN ACEDIVRC ACEHURT1
94 ACEPRISN ACEDIVRC ACESWEAR
95 ACEPRISN ACEDIVRC ACESEX
96 ACEPRISN ACEDIVRC ACEADSAF
97 ACEPRISN ACEDIVRC ACEADNED
98 ACEPRISN ACEDIVRC ACEDEPRS
99 ACEPRISN ACEPUNCH ACEHURT1
100 ACEPRISN ACEPUNCH ACESWEAR
101 ACEPRISN ACEPUNCH ACESEX
102 ACEPRISN ACEPUNCH ACEADSAF
103 ACEPRISN ACEPUNCH ACEADNED
104 ACEPRISN ACEPUNCH ACEDEPRS
105 ACEPRISN ACEHURT1 ACESWEAR
106 ACEPRISN ACEHURT1 ACESEX
107 ACEPRISN ACEHURT1 ACEADSAF
108 ACEPRISN ACEHURT1 ACEADNED
109 ACEPRISN ACEHURT1 ACEDEPRS
110 ACEPRISN ACESWEAR ACESEX
111 ACEPRISN ACESWEAR ACEADSAF
112 ACEPRISN ACESWEAR ACEADNED
113 ACEPRISN ACESWEAR ACEDEPRS
114 ACEPRISN ACESEX ACEADSAF
115 ACEPRISN ACESEX ACEADNED
116 ACEPRISN ACESEX ACEDEPRS
117 ACEPRISN ACEADSAF ACEADNED
118 ACEPRISN ACEADSAF ACEDEPRS
119 ACEPRISN ACEADNED ACEDEPRS
120 ACEDIVRC ACEPUNCH ACEHURT1
121 ACEDIVRC ACEPUNCH ACESWEAR
122 ACEDIVRC ACEPUNCH ACESEX
123 ACEDIVRC ACEPUNCH ACEADSAF
124 ACEDIVRC ACEPUNCH ACEADNED
125 ACEDIVRC ACEPUNCH ACEDEPRS
126 ACEDIVRC ACEHURT1 ACESWEAR
127 ACEDIVRC ACEHURT1 ACESEX
128 ACEDIVRC ACEHURT1 ACEADSAF
129 ACEDIVRC ACEHURT1 ACEADNED
130 ACEDIVRC ACEHURT1 ACEDEPRS
131 ACEDIVRC ACESWEAR ACESEX
132 ACEDIVRC ACESWEAR ACEADSAF
133 ACEDIVRC ACESWEAR ACEADNED
134 ACEDIVRC ACESWEAR ACEDEPRS
135 ACEDIVRC ACESEX ACEADSAF
136 ACEDIVRC ACESEX ACEADNED
137 ACEDIVRC ACESEX ACEDEPRS
138 ACEDIVRC ACEADSAF ACEADNED
139 ACEDIVRC ACEADSAF ACEDEPRS
140 ACEDIVRC ACEADNED ACEDEPRS
141 ACEPUNCH ACEHURT1 ACESWEAR
142 ACEPUNCH ACEHURT1 ACESEX
143 ACEPUNCH ACEHURT1 ACEADSAF
144 ACEPUNCH ACEHURT1 ACEADNED
145 ACEPUNCH ACEHURT1 ACEDEPRS
146 ACEPUNCH ACESWEAR ACESEX
147 ACEPUNCH ACESWEAR ACEADSAF
148 ACEPUNCH ACESWEAR ACEADNED
149 ACEPUNCH ACESWEAR ACEDEPRS
150 ACEPUNCH ACESEX ACEADSAF
151 ACEPUNCH ACESEX ACEADNED
152 ACEPUNCH ACESEX ACEDEPRS
153 ACEPUNCH ACEADSAF ACEADNED
154 ACEPUNCH ACEADSAF ACEDEPRS
155 ACEPUNCH ACEADNED ACEDEPRS
156 ACEHURT1 ACESWEAR ACESEX
157 ACEHURT1 ACESWEAR ACEADSAF
158 ACEHURT1 ACESWEAR ACEADNED
159 ACEHURT1 ACESWEAR ACEDEPRS
160 ACEHURT1 ACESEX ACEADSAF
161 ACEHURT1 ACESEX ACEADNED
162 ACEHURT1 ACESEX ACEDEPRS
163 ACEHURT1 ACEADSAF ACEADNED
164 ACEHURT1 ACEADSAF ACEDEPRS
165 ACEHURT1 ACEADNED ACEDEPRS
166 ACESWEAR ACESEX ACEADSAF
167 ACESWEAR ACESEX ACEADNED
168 ACESWEAR ACESEX ACEDEPRS
169 ACESWEAR ACEADSAF ACEADNED
170 ACESWEAR ACEADSAF ACEDEPRS
171 ACESWEAR ACEADNED ACEDEPRS
172 ACESEX ACEADSAF ACEADNED
173 ACESEX ACEADSAF ACEDEPRS
174 ACESEX ACEADNED ACEDEPRS
175 ACEADSAF ACEADNED ACEDEPRS
176 ACESUB ACEPRISN ACEDIVRC ACEPUNCH
177 ACESUB ACEPRISN ACEDIVRC ACEHURT1
178 ACESUB ACEPRISN ACEDIVRC ACESWEAR
179 ACESUB ACEPRISN ACEDIVRC ACESEX
180 ACESUB ACEPRISN ACEDIVRC ACEADSAF
181 ACESUB ACEPRISN ACEDIVRC ACEADNED
182 ACESUB ACEPRISN ACEDIVRC ACEDEPRS
183 ACESUB ACEPRISN ACEPUNCH ACEHURT1
184 ACESUB ACEPRISN ACEPUNCH ACESWEAR
185 ACESUB ACEPRISN ACEPUNCH ACESEX
186 ACESUB ACEPRISN ACEPUNCH ACEADSAF
187 ACESUB ACEPRISN ACEPUNCH ACEADNED
188 ACESUB ACEPRISN ACEPUNCH ACEDEPRS
189 ACESUB ACEPRISN ACEHURT1 ACESWEAR
190 ACESUB ACEPRISN ACEHURT1 ACESEX
191 ACESUB ACEPRISN ACEHURT1 ACEADSAF
192 ACESUB ACEPRISN ACEHURT1 ACEADNED
193 ACESUB ACEPRISN ACEHURT1 ACEDEPRS
194 ACESUB ACEPRISN ACESWEAR ACESEX
195 ACESUB ACEPRISN ACESWEAR ACEADSAF
196 ACESUB ACEPRISN ACESWEAR ACEADNED
197 ACESUB ACEPRISN ACESWEAR ACEDEPRS
198 ACESUB ACEPRISN ACESEX ACEADSAF
199 ACESUB ACEPRISN ACESEX ACEADNED
200 ACESUB ACEPRISN ACESEX ACEDEPRS
201 ACESUB ACEPRISN ACEADSAF ACEADNED
202 ACESUB ACEPRISN ACEADSAF ACEDEPRS
203 ACESUB ACEPRISN ACEADNED ACEDEPRS
204 ACESUB ACEDIVRC ACEPUNCH ACEHURT1
205 ACESUB ACEDIVRC ACEPUNCH ACESWEAR
206 ACESUB ACEDIVRC ACEPUNCH ACESEX
207 ACESUB ACEDIVRC ACEPUNCH ACEADSAF
208 ACESUB ACEDIVRC ACEPUNCH ACEADNED
209 ACESUB ACEDIVRC ACEPUNCH ACEDEPRS
210 ACESUB ACEDIVRC ACEHURT1 ACESWEAR
211 ACESUB ACEDIVRC ACEHURT1 ACESEX
212 ACESUB ACEDIVRC ACEHURT1 ACEADSAF
213 ACESUB ACEDIVRC ACEHURT1 ACEADNED
214 ACESUB ACEDIVRC ACEHURT1 ACEDEPRS
215 ACESUB ACEDIVRC ACESWEAR ACESEX
216 ACESUB ACEDIVRC ACESWEAR ACEADSAF
217 ACESUB ACEDIVRC ACESWEAR ACEADNED
218 ACESUB ACEDIVRC ACESWEAR ACEDEPRS
219 ACESUB ACEDIVRC ACESEX ACEADSAF
220 ACESUB ACEDIVRC ACESEX ACEADNED
221 ACESUB ACEDIVRC ACESEX ACEDEPRS
222 ACESUB ACEDIVRC ACEADSAF ACEADNED
223 ACESUB ACEDIVRC ACEADSAF ACEDEPRS
224 ACESUB ACEDIVRC ACEADNED ACEDEPRS
225 ACESUB ACEPUNCH ACEHURT1 ACESWEAR
226 ACESUB ACEPUNCH ACEHURT1 ACESEX
227 ACESUB ACEPUNCH ACEHURT1 ACEADSAF
228 ACESUB ACEPUNCH ACEHURT1 ACEADNED
229 ACESUB ACEPUNCH ACEHURT1 ACEDEPRS
230 ACESUB ACEPUNCH ACESWEAR ACESEX
231 ACESUB ACEPUNCH ACESWEAR ACEADSAF
232 ACESUB ACEPUNCH ACESWEAR ACEADNED
233 ACESUB ACEPUNCH ACESWEAR ACEDEPRS
234 ACESUB ACEPUNCH ACESEX ACEADSAF
235 ACESUB ACEPUNCH ACESEX ACEADNED
236 ACESUB ACEPUNCH ACESEX ACEDEPRS
237 ACESUB ACEPUNCH ACEADSAF ACEADNED
238 ACESUB ACEPUNCH ACEADSAF ACEDEPRS
239 ACESUB ACEPUNCH ACEADNED ACEDEPRS
240 ACESUB ACEHURT1 ACESWEAR ACESEX
241 ACESUB ACEHURT1 ACESWEAR ACEADSAF
242 ACESUB ACEHURT1 ACESWEAR ACEADNED
243 ACESUB ACEHURT1 ACESWEAR ACEDEPRS
244 ACESUB ACEHURT1 ACESEX ACEADSAF
245 ACESUB ACEHURT1 ACESEX ACEADNED
246 ACESUB ACEHURT1 ACESEX ACEDEPRS
247 ACESUB ACEHURT1 ACEADSAF ACEADNED
248 ACESUB ACEHURT1 ACEADSAF ACEDEPRS
249 ACESUB ACEHURT1 ACEADNED ACEDEPRS
250 ACESUB ACESWEAR ACESEX ACEADSAF
251 ACESUB ACESWEAR ACESEX ACEADNED
252 ACESUB ACESWEAR ACESEX ACEDEPRS
253 ACESUB ACESWEAR ACEADSAF ACEADNED
254 ACESUB ACESWEAR ACEADSAF ACEDEPRS
255 ACESUB ACESWEAR ACEADNED ACEDEPRS
256 ACESUB ACESEX ACEADSAF ACEADNED
257 ACESUB ACESEX ACEADSAF ACEDEPRS
258 ACESUB ACESEX ACEADNED ACEDEPRS
259 ACESUB ACEADSAF ACEADNED ACEDEPRS
260 ACEPRISN ACEDIVRC ACEPUNCH ACEHURT1
261 ACEPRISN ACEDIVRC ACEPUNCH ACESWEAR
262 ACEPRISN ACEDIVRC ACEPUNCH ACESEX
263 ACEPRISN ACEDIVRC ACEPUNCH ACEADSAF
264 ACEPRISN ACEDIVRC ACEPUNCH ACEADNED
265 ACEPRISN ACEDIVRC ACEPUNCH ACEDEPRS
266 ACEPRISN ACEDIVRC ACEHURT1 ACESWEAR
267 ACEPRISN ACEDIVRC ACEHURT1 ACESEX
268 ACEPRISN ACEDIVRC ACEHURT1 ACEADSAF
269 ACEPRISN ACEDIVRC ACEHURT1 ACEADNED
270 ACEPRISN ACEDIVRC ACEHURT1 ACEDEPRS
271 ACEPRISN ACEDIVRC ACESWEAR ACESEX
272 ACEPRISN ACEDIVRC ACESWEAR ACEADSAF
273 ACEPRISN ACEDIVRC ACESWEAR ACEADNED
274 ACEPRISN ACEDIVRC ACESWEAR ACEDEPRS
275 ACEPRISN ACEDIVRC ACESEX ACEADSAF
276 ACEPRISN ACEDIVRC ACESEX ACEADNED
277 ACEPRISN ACEDIVRC ACESEX ACEDEPRS
278 ACEPRISN ACEDIVRC ACEADSAF ACEADNED
279 ACEPRISN ACEDIVRC ACEADSAF ACEDEPRS
280 ACEPRISN ACEDIVRC ACEADNED ACEDEPRS
281 ACEPRISN ACEPUNCH ACEHURT1 ACESWEAR
282 ACEPRISN ACEPUNCH ACEHURT1 ACESEX
283 ACEPRISN ACEPUNCH ACEHURT1 ACEADSAF
284 ACEPRISN ACEPUNCH ACEHURT1 ACEADNED
285 ACEPRISN ACEPUNCH ACEHURT1 ACEDEPRS
286 ACEPRISN ACEPUNCH ACESWEAR ACESEX
287 ACEPRISN ACEPUNCH ACESWEAR ACEADSAF
288 ACEPRISN ACEPUNCH ACESWEAR ACEADNED
289 ACEPRISN ACEPUNCH ACESWEAR ACEDEPRS
290 ACEPRISN ACEPUNCH ACESEX ACEADSAF
291 ACEPRISN ACEPUNCH ACESEX ACEADNED
292 ACEPRISN ACEPUNCH ACESEX ACEDEPRS
293 ACEPRISN ACEPUNCH ACEADSAF ACEADNED
294 ACEPRISN ACEPUNCH ACEADSAF ACEDEPRS
295 ACEPRISN ACEPUNCH ACEADNED ACEDEPRS
296 ACEPRISN ACEHURT1 ACESWEAR ACESEX
297 ACEPRISN ACEHURT1 ACESWEAR ACEADSAF
298 ACEPRISN ACEHURT1 ACESWEAR ACEADNED
299 ACEPRISN ACEHURT1 ACESWEAR ACEDEPRS
300 ACEPRISN ACEHURT1 ACESEX ACEADSAF
301 ACEPRISN ACEHURT1 ACESEX ACEADNED
302 ACEPRISN ACEHURT1 ACESEX ACEDEPRS
303 ACEPRISN ACEHURT1 ACEADSAF ACEADNED
304 ACEPRISN ACEHURT1 ACEADSAF ACEDEPRS
305 ACEPRISN ACEHURT1 ACEADNED ACEDEPRS
306 ACEPRISN ACESWEAR ACESEX ACEADSAF
307 ACEPRISN ACESWEAR ACESEX ACEADNED
308 ACEPRISN ACESWEAR ACESEX ACEDEPRS
309 ACEPRISN ACESWEAR ACEADSAF ACEADNED
310 ACEPRISN ACESWEAR ACEADSAF ACEDEPRS
311 ACEPRISN ACESWEAR ACEADNED ACEDEPRS
312 ACEPRISN ACESEX ACEADSAF ACEADNED
313 ACEPRISN ACESEX ACEADSAF ACEDEPRS
314 ACEPRISN ACESEX ACEADNED ACEDEPRS
315 ACEPRISN ACEADSAF ACEADNED ACEDEPRS
316 ACEDIVRC ACEPUNCH ACEHURT1 ACESWEAR
317 ACEDIVRC ACEPUNCH ACEHURT1 ACESEX
318 ACEDIVRC ACEPUNCH ACEHURT1 ACEADSAF
319 ACEDIVRC ACEPUNCH ACEHURT1 ACEADNED
320 ACEDIVRC ACEPUNCH ACEHURT1 ACEDEPRS
321 ACEDIVRC ACEPUNCH ACESWEAR ACESEX
322 ACEDIVRC ACEPUNCH ACESWEAR ACEADSAF
323 ACEDIVRC ACEPUNCH ACESWEAR ACEADNED
324 ACEDIVRC ACEPUNCH ACESWEAR ACEDEPRS
325 ACEDIVRC ACEPUNCH ACESEX ACEADSAF
326 ACEDIVRC ACEPUNCH ACESEX ACEADNED
327 ACEDIVRC ACEPUNCH ACESEX ACEDEPRS
328 ACEDIVRC ACEPUNCH ACEADSAF ACEADNED
329 ACEDIVRC ACEPUNCH ACEADSAF ACEDEPRS
330 ACEDIVRC ACEPUNCH ACEADNED ACEDEPRS
331 ACEDIVRC ACEHURT1 ACESWEAR ACESEX
332 ACEDIVRC ACEHURT1 ACESWEAR ACEADSAF
333 ACEDIVRC ACEHURT1 ACESWEAR ACEADNED
334 ACEDIVRC ACEHURT1 ACESWEAR ACEDEPRS
335 ACEDIVRC ACEHURT1 ACESEX ACEADSAF
336 ACEDIVRC ACEHURT1 ACESEX ACEADNED
337 ACEDIVRC ACEHURT1 ACESEX ACEDEPRS
338 ACEDIVRC ACEHURT1 ACEADSAF ACEADNED
339 ACEDIVRC ACEHURT1 ACEADSAF ACEDEPRS
340 ACEDIVRC ACEHURT1 ACEADNED ACEDEPRS
341 ACEDIVRC ACESWEAR ACESEX ACEADSAF
342 ACEDIVRC ACESWEAR ACESEX ACEADNED
343 ACEDIVRC ACESWEAR ACESEX ACEDEPRS
344 ACEDIVRC ACESWEAR ACEADSAF ACEADNED
345 ACEDIVRC ACESWEAR ACEADSAF ACEDEPRS
346 ACEDIVRC ACESWEAR ACEADNED ACEDEPRS
347 ACEDIVRC ACESEX ACEADSAF ACEADNED
348 ACEDIVRC ACESEX ACEADSAF ACEDEPRS
349 ACEDIVRC ACESEX ACEADNED ACEDEPRS
350 ACEDIVRC ACEADSAF ACEADNED ACEDEPRS
351 ACEPUNCH ACEHURT1 ACESWEAR ACESEX
352 ACEPUNCH ACEHURT1 ACESWEAR ACEADSAF
353 ACEPUNCH ACEHURT1 ACESWEAR ACEADNED
354 ACEPUNCH ACEHURT1 ACESWEAR ACEDEPRS
355 ACEPUNCH ACEHURT1 ACESEX ACEADSAF
356 ACEPUNCH ACEHURT1 ACESEX ACEADNED
357 ACEPUNCH ACEHURT1 ACESEX ACEDEPRS
358 ACEPUNCH ACEHURT1 ACEADSAF ACEADNED
359 ACEPUNCH ACEHURT1 ACEADSAF ACEDEPRS
360 ACEPUNCH ACEHURT1 ACEADNED ACEDEPRS
361 ACEPUNCH ACESWEAR ACESEX ACEADSAF
362 ACEPUNCH ACESWEAR ACESEX ACEADNED
363 ACEPUNCH ACESWEAR ACESEX ACEDEPRS
364 ACEPUNCH ACESWEAR ACEADSAF ACEADNED
365 ACEPUNCH ACESWEAR ACEADSAF ACEDEPRS
366 ACEPUNCH ACESWEAR ACEADNED ACEDEPRS
367 ACEPUNCH ACESEX ACEADSAF ACEADNED
368 ACEPUNCH ACESEX ACEADSAF ACEDEPRS
369 ACEPUNCH ACESEX ACEADNED ACEDEPRS
370 ACEPUNCH ACEADSAF ACEADNED ACEDEPRS
371 ACEHURT1 ACESWEAR ACESEX ACEADSAF
372 ACEHURT1 ACESWEAR ACESEX ACEADNED
373 ACEHURT1 ACESWEAR ACESEX ACEDEPRS
374 ACEHURT1 ACESWEAR ACEADSAF ACEADNED
375 ACEHURT1 ACESWEAR ACEADSAF ACEDEPRS
376 ACEHURT1 ACESWEAR ACEADNED ACEDEPRS
377 ACEHURT1 ACESEX ACEADSAF ACEADNED
378 ACEHURT1 ACESEX ACEADSAF ACEDEPRS
379 ACEHURT1 ACESEX ACEADNED ACEDEPRS
380 ACEHURT1 ACEADSAF ACEADNED ACEDEPRS
381 ACESWEAR ACESEX ACEADSAF ACEADNED
382 ACESWEAR ACESEX ACEADSAF ACEDEPRS
383 ACESWEAR ACESEX ACEADNED ACEDEPRS
384 ACESWEAR ACEADSAF ACEADNED ACEDEPRS
385 ACESEX ACEADSAF ACEADNED ACEDEPRS
Appendix B Sample Sizes by Stratum in Each Subgroup
Table 4. Sample Sizes by Stratum in the Blue Subgroup
Race/Ethnicity Sex Age Sample Size
1 White, Non-Hispanic Male 18-24 362
2 Black, Non-Hispanic Male 18-24 61
3 Hispanic Male 18-24 182
4 Other race, Non-Hispanic Male 18-24 91
5 White, Non-Hispanic Female 18-24 232
6 Black, Non-Hispanic Female 18-24 63
7 Hispanic Female 18-24 135
8 Other race, Non-Hispanic Female 18-24 72
9 White, Non-Hispanic Male 25-34 695
10 Black, Non-Hispanic Male 25-34 82
11 Hispanic Male 25-34 234
12 Other race, Non-Hispanic Male 25-34 133
13 White, Non-Hispanic Female 25-34 621
14 Black, Non-Hispanic Female 25-34 99
15 Hispanic Female 25-34 274
16 Other race, Non-Hispanic Female 25-34 125
17 White, Non-Hispanic Male 35-44 820
18 Black, Non-Hispanic Male 35-44 103
19 Hispanic Male 35-44 222
20 Other race, Non-Hispanic Male 35-44 194
21 White, Non-Hispanic Female 35-44 908
22 Black, Non-Hispanic Female 35-44 157
23 Hispanic Female 35-44 247
24 Other race, Non-Hispanic Female 35-44 129
25 White, Non-Hispanic Male 45-54 955
26 Black, Non-Hispanic Male 45-54 135
27 Hispanic Male 45-54 171
28 Other race, Non-Hispanic Male 45-54 131
29 White, Non-Hispanic Female 45-54 1212
30 Black, Non-Hispanic Female 45-54 168
31 Hispanic Female 45-54 206
32 Other race, Non-Hispanic Female 45-54 151
33 White, Non-Hispanic Male 55-64 1513
34 Black, Non-Hispanic Male 55-64 141
35 Hispanic Male 55-64 160
36 Other race, Non-Hispanic Male 55-64 132
37 White, Non-Hispanic Female 55-64 2054
38 Black, Non-Hispanic Female 55-64 267
39 Hispanic Female 55-64 171
40 Other race, Non-Hispanic Female 55-64 110
41 White, Non-Hispanic Male 65+ 3688
42 Black, Non-Hispanic Male 65+ 260
43 Hispanic Male 65+ 118
44 Other race, Non-Hispanic Male 65+ 180
45 White, Non-Hispanic Female 65+ 4561
46 Black, Non-Hispanic Female 65+ 473
47 Hispanic Female 65+ 139
48 Other race, Non-Hispanic Female 65+ 169
Table 5. Sample Sizes by Stratum in the Red Subgroup
Race/Ethnicity Sex Age Sample Size
1 White, Non-Hispanic Male 18-24 362
2 Black, Non-Hispanic Male 18-24 61
3 Hispanic Male 18-24 182
4 Other race, Non-Hispanic Male 18-24 91
5 White, Non-Hispanic Female 18-24 232
6 Black, Non-Hispanic Female 18-24 63
7 Hispanic Female 18-24 135
8 Other race, Non-Hispanic Female 18-24 72
9 White, Non-Hispanic Male 25-34 695
10 Black, Non-Hispanic Male 25-34 82
11 Hispanic Male 25-34 234
12 Other race, Non-Hispanic Male 25-34 133
13 White, Non-Hispanic Female 25-34 621
14 Black, Non-Hispanic Female 25-34 99
15 Hispanic Female 25-34 274
16 Other race, Non-Hispanic Female 25-34 125
17 White, Non-Hispanic Male 35-44 820
18 Black, Non-Hispanic Male 35-44 103
19 Hispanic Male 35-44 222
20 Other race, Non-Hispanic Male 35-44 194
21 White, Non-Hispanic Female 35-44 908
22 Black, Non-Hispanic Female 35-44 157
23 Hispanic Female 35-44 247
24 Other race, Non-Hispanic Female 35-44 129
25 White, Non-Hispanic Male 45-54 955
26 Black, Non-Hispanic Male 45-54 135
27 Hispanic Male 45-54 171
28 Other race, Non-Hispanic Male 45-54 131
29 White, Non-Hispanic Female 45-54 1212
30 Black, Non-Hispanic Female 45-54 168
31 Hispanic Female 45-54 206
32 Other race, Non-Hispanic Female 45-54 151
33 White, Non-Hispanic Male 55-64 1513
34 Black, Non-Hispanic Male 55-64 141
35 Hispanic Male 55-64 160
36 Other race, Non-Hispanic Male 55-64 132
37 White, Non-Hispanic Female 55-64 2054
38 Black, Non-Hispanic Female 55-64 267
39 Hispanic Female 55-64 171
40 Other race, Non-Hispanic Female 55-64 110
41 White, Non-Hispanic Male 65+ 3688
42 Black, Non-Hispanic Male 65+ 260
43 Hispanic Male 65+ 118
44 Other race, Non-Hispanic Male 65+ 180
45 White, Non-Hispanic Female 65+ 4561
46 Black, Non-Hispanic Female 65+ 473
47 Hispanic Female 65+ 139
48 Other race, Non-Hispanic Female 65+ 169

### v2

@@
 Although these can take various forms, they are typically categorized into ten types involving abuse (physical, emotional, or sexual) and neglect (physical or emotional), as well as household challenges including mental illness, substance abuse, intimate partner violence, parental absence due to divorce or separation, or incarceration.
-Initially investigated in a landmark study conducted by the Centers for Disease Control and Prevention (CDC) and Kaiser Permanente, ACEs are highly prevalent and exhibit a strong dose-response relationship with numerous negative health and social outcomes in adulthood (Felitti et al.
-[1998]; Dube et al.
-[2003]; Anda et al.
-[2006]).
+ACEs were initially investigated in a landmark study conducted by the Centers for Disease Control and Prevention (CDC) and Kaiser Permanente, and are highly prevalent and exhibit a strong dose-response relationship with numerous negative health and social outcomes in adulthood (felitti1998relationship; dube2003impact; anda2006enduring).
 Exposure to individual ACEs has been repeatedly associated with a range of adverse physical and emotional health consequences later in life.
-One study indicated that childhood emotional abuse increases the risk of lifetime depressive disorders, with adjusted odds ratios (ORs) of 2.7 for women and 2.5 for men (Chapman et al.
-[2004]).
-Depression also appears to be closely related to having experienced parental alcohol abuse (Anda et al.
-[2002]).
-Moreover, physical and verbal abuse have been found to be strongly associated with body weight issues and obesity (Williamson et al.
-[2002]).
-Unfortunately, when children are exposed to one type of adversity, there is a high likelihood of exposure to additional ACEs (Dong et al.
-[2004]; Felitti [2002]).
-Numerous large-scale investigations have confirmed that these detrimental experiences can have a cumulative effect: each additional type of adversity elevates the probability of unfavorable outcomes (Felitti et al.
-[1998]; Chapman et al.
-[2004]).
-Accordingly, previous literature has examined closely the number of categories of ACEs, often referred to as ACE score.
-For instance, one study using data from 14 states revealed substantial variation across states in the percentage change in health outcomes for those exposed to at least four ACEs compared to those with fewer than four ACEs (Waehrer et al.
-[2020]).
-Physically, individuals with four or more ACEs face elevated risks of heart disease, liver disease, kidney disease, asthma, Alzheimer’s disease, stroke, metabolic disorders, obesity, diabetes, and certain cancers (for Disease Control and Prevention [2017]; for Youth Wellness [2014]; Merrick [2019]; Hughes et al.
-[2017]; Waehrer et al.
-[2020]).
-On the emotional and psychological side, ACE score strongly correlates with mental health issues such as depression, anxiety, and suicidal ideation (Hillis et al.
-[2004]; Chapman et al.
-[2004]).
-Meta-analyses further reinforce that an accumulation of at least 4 ACEs correlates with higher risks of chronic disease, poor mental health, and negative social outcomes, demonstrating the need for strategies to prevent ACEs and lessen their long-term repercussions (Hughes et al.
-[2017]).
-A recent study challenged the overall ACE score and advised to break down the measure back into specific items (Fitzgerald and Bishop [2024]).
+One study has found that childhood emotional abuse increases the risk of lifetime depressive disorders, with adjusted odds ratios (ORs) of 2.7 for women and 2.5 for men (chapman2004adverse).
+Depression also appears to be closely related to having experienced parental alcohol abuse (anda2002adverse).
+Moreover, physical and verbal abuse have been found to be strongly associated with body weight issues and obesity (williamson2002body).
+Unfortunately, when children are exposed to one type of adversity, there is a high likelihood of exposure to additional ACEs (dong2004interrelatedness; felitti2002relationship).
+Numerous large-scale investigations have confirmed that these detrimental experiences can have a cumulative effect: each additional type of adversity elevates the probability of unfavorable outcomes (felitti1998relationship; chapman2004adverse).
+Accordingly, previous literature has closely examined the number of categories of ACEs, often referred to as the ACE score.
+For instance, one study using data from 14 states revealed substantial variation across states in the percentage change in health outcomes for those exposed to at least four ACEs compared to those with fewer than four ACEs (waehrer2020disease).
+Physically, individuals with four or more ACEs face elevated risks of heart disease, liver disease, kidney disease, asthma, Alzheimer’s disease, stroke, metabolic disorders, obesity, diabetes, and certain cancers (CDC2017; CYW2014; merrick2019vital; hughes2017effect; waehrer2020disease).
+On the emotional and psychological side, the ACE score strongly correlates with mental health issues such as depression, anxiety, and suicidal ideation (hillis2004association; chapman2004adverse).
+Meta-analyses further reinforce that an accumulation of at least 4 ACEs correlates with higher risks of chronic disease, poor mental health, and negative social outcomes, demonstrating the need for strategies to prevent ACEs and lessen their long-term repercussions (hughes2017effect).
+A growing literature cautions that the additive ACE score is a blunt tool for individual risk, as it treats all events as equal.
+Then, a child enduring repeated sexual abuse is assumed the same as exposure to parental divorce, which appears problematic on its face.
+The ACE score can encourage over-simplified claims about risk or causality, deterministic framing, and stigma (lacey2020practitioner).
+Methodologically, large cohort analyses show that an ACE score cutoff such as $\geq 4$ poorly distinguishes which specific individuals will develop later health problems (baldwin2021population; meehan2022poor).
+A cohort study of women in correctional custody showed that maltreatment alone is significantly associated with adult mental health while household dysfunction is not.
+Therefore, the authors challenged the overall ACE score and recommended decomposing the measure back into specific items (fitzgerald2024challenging).
+Recent pediatric guidance finds insufficient evidence to support routine, threshold-based ACE screening in clinical settings and supports trauma-informed, comprehensive assessment instead (austin2024screening).
 Gaps also remain in our understanding of how the accumulation of ACEs influences health.
-We are therefore interested in which categories of ACEs—and which specific combinations of ACEs (e.g., physical abuse + sexual abuse, or physical abuse + emotional neglect + substance abuse)—are most harmful and warrant primary care.
+We are therefore interested in which categories of ACEs—and which specific combinations of ACEs (e.g., physical abuse + sexual abuse, or physical abuse + emotional neglect + substance abuse)—are most harmful and warrant particular attention in primary care settings.
 We focus our analysis on depression, one of the most prevalent and burdensome mental health conditions globally.
-Consequently, our focus is to evaluate the effect of different combinations of specific ACE items on depression and identifying a higher-risk group of sets of ACEs.
+Consequently, our focus is to evaluate the effect of different combinations of specific ACE items on depression and to identify a higher-risk group of sets of ACEs.
 We will examine the impacts of combinations of ACEs on adult depression using data from the Behavioral Risk Factor Surveillance System (BRFSS) 2023, an ongoing health-related telephone survey system administered by the CDC with a random sample of over 400,000 individuals across the United States in 2023.
-This data includes detailed information from ten states that administered the CDC’s Adverse Childhood Experiences module in 2023.
-We can obtain all ten ACE categories proposed in the CDC-Kaiser study (Felitti et al.
-[1998]), and there are over 50,000 individuals with complete records of all ACE items.
+These data include detailed information from ten states that administered the CDC’s Adverse Childhood Experiences module in 2023.
+We observe all ten ACE categories proposed in the CDC-Kaiser study (felitti1998relationship), and there are over 50,000 individuals with complete records of all ACE items.
 Among them, around 17,000 reported zero ACEs while the rest were exposed to at least one ACE.
-This large sample size will help facilitate powerful data analysis and explore the influencing mechanism of different ACEs combinations.
-The BRFSS also collected measurements of variables that may confound the relationship between exposure to ACEs and depression, such as those related to personal demographics.
-We will utilize those variables to form separate strata for treated and control individuals, enabling us to compare the outcomes of similar individuals in hopes of ameliorating confounding biases.
+This large sample size will help support powerful data analysis and facilitate the exploration of influencing mechanism of different ACEs combinations.
 We employ a novel statistical design called data turnover (Bekerman et al.
-in preparation, see Bekerman et al.
-[2024a] for a description) to conduct this analysis.
+in preparation, see bekerman2024protocol for a description) to conduct this analysis.
 By integrating both qualitative and quantitative insights from data exploration, data turnover enables a single team of statisticians and domain experts to evaluate evidence from multiple data splits.
 Exploratory data analysis (EDA) allows us to inspect the data before making potentially questionable assumptions, detect anomalies or outliers that do not represent natural variation in the data, and formulate new hypotheses.
-It also helps identify variables that may not measure what we initially intended, indicating the need for a better measure or refined, data-informed hypotheses (Bekerman et al.
-[2024b]).
+It also helps identify variables that may not measure what we initially intended, indicating the need for a better measure or refined, data-informed hypotheses (bekerman2024planning).
 Beyond exploration, replication is critical in observational studies for reinforcing the credibility of findings and sustaining trust in scientific knowledge.
 Previous work has combined some—but not all—aspects of exploration and replication within the same observational study.
-For instance, some methods allow for exploration without replication (e.g., Cox [1975]; Heller et al.
-[2009]), whereas others enable replication without exploration (e.g., Zhao et al.
-[2018]; Karmakar et al.
-[2019]).
-There is also research that supports both exploration and replication but requires two independent investigative teams (Roy et al.
-[2022]).
+For instance, some methods allow for exploration without replication (e.g., cox1975note; heller2009split), whereas others enable replication without exploration (e.g., zhao2018cross; karmakar2019integrating).
+There is also research that supports both exploration and replication but requires two independent investigative teams (roy2022protocol).
 In contrast, data turnover makes it possible to conduct both exploration and replication using a single research group.
-In observational studies like ours, where treatment assignment is not under our direct control, obtaining consistent findings across multiple groups that may have different assignment mechanisms reinforces the evidence that the treatment itself is truly responsible for the observed effects (Rosenbaum [2001]; Rosenbaum [2015]).
+In observational studies like ours, where treatment assignment is not under our direct control, obtaining consistent findings across multiple groups that may have different assignment mechanisms reinforces the evidence that the treatment itself is truly responsible for the observed effects (rosenbaum2001replicating; rosenbaum2015see).
 By analyzing data from two distinct subpopulations of BRFSS 2023 participants—residing in states that voted for the Democratic candidate in the 2024 US presidential election (”blue states”) and those residing in states that voted for the Republican candidate (”red states”) —we can assess the replicability of our findings.
 We split our data into blue states and red states, recognizing that these political orientations often shape distinct social, economic, and health care policies.
-A study focusing on population in East and West Germany has shown that socio-political context is a determinant of childhood maltreatment (Ulke et al.
-[2021]).
+A study focusing on the population in East and West Germany has shown that socio-political context is a determinant of childhood maltreatment (ulke2021socio).
 In our context, for example, blue states may allocate more resources toward public health programs (e.g., adopt Medicaid expansion) or have broader social safety nets, whereas red states may emphasize different strategies for healthcare funding and child protective services.
@@
 By employing data turnover, we can predetermine how one subgroup will guide analysis of the other, then apply exploratory data analysis to the second subgroup to inform how we draw inferences about the first.
-Any outcome that shows a significant effect in both groups is defined as a “replicable” finding.
-(Bogomolov and Heller [2023]).
+Any outcome that shows a significant effect in both groups is defined as a “replicable” finding (bogomolov2023replicability).
 In Section 4, we detail how we apply this non-random cross-screening method to detect replicable outcomes and test the global null hypothesis, i.e., identifying outcomes that are affected in at least one of the two subpopulations.
@@
 We introduce the BRFSS dataset and describe our data processing in Section 2.
-In Section 3, we report our candidate sets of combinations.
-In Section 4, we outline our stratification methodology.
-Finally, we conclude with a description of our data analysis procedure and explanation of several specifics of implementation in Section 5.
+In Section 3, we outline our methodology.
+Finally, we conclude with a description of our data analysis procedure and an explanation of several specifics of implementation in Section 4.
 2. The Behavioral Risk Factor Surveillance System (BRFSS)
-This study utilizes data from BRFSS, an ongoing health-related telephone survey system administered by the CDC.
+This study uses data from BRFSS, an ongoing health-related telephone survey system administered by the CDC.
 BRFSS collects comprehensive data on a wide range of health-related behaviors, chronic health conditions, healthcare access, and preventive service usage from the non-institutionalized adult population (18 years and older).
@@
 sample that administered the CDC’s ACEs module in 2023.
-Following previous studies utilizing the BRFSS data (Bhan et al.
-[2014]; Gilbert et al.
-[2015]; Waehrer et al.
-[2020]), we collapse different questions into particular indicators.
+Following previous studies utilizing the BRFSS data (bhan2014childhood; gilbert2015childhood; waehrer2020disease), we collapse different questions into binary indicators of each ACE.
 More details are provided in Table 1.
-In total, we create 10 ACEs items from the BRFSS data (see table 2), which is consistent with the categorization in CDC-Kaiser study (Felitti et al.
-[1998]).
-It is worth noting that we first follow those criteria for data pre-processing, but these steps could be adjusted when we carry out exploratory data analysis in the data turnover..
-Indicator Questions Collapsed Encoding Method
-Any Sexual Abuse Number of times forced to touch, be touched, or have sex with anyone at least five years older or an adult Collapsed into a single indicator for any sexual abuse
-Household Substance Use Alcohol and illegal drug use in the household Combined into a single indicator for household substance use
-Physical Abuse Frequency of physical abuse Encoded as 1 if the answer is “once” or “more than once”
-Verbal Abuse Frequency of verbal abuse Encoded as 1 if the answer is “once” or “more than once”
-Parental Divorce Response to parental divorce question Affirmative response encoded as 1
-Living with a Depressed or Mentally Ill Person Response to question about living with a mentally ill person Affirmative response encoded as 1
-Living with an Incarcerated Person Response to question about living with an incarcerated person Affirmative response encoded as 1
-Physical Violence Between Parents Frequency of physical violence between parents Affirmative response encoded as 1
-Safety Did an adult make you feel safe and protected ‘Never’ and ‘A little of the time’ encoded as 0, ‘Some of the time’, ‘Most of the time’, ‘All of the time’ encoded as 1
-Basic Needs Met Did an adult make sure basic needs were met ‘Never’ and ‘A little of the time’ encoded as 0, ‘Some of the time’, ‘Most of the time’, ‘All of the time’ encoded as 1
+In total, we create 10 binary ACE items from the BRFSS data (see Table 2), which is consistent with the categorization in the CDC-Kaiser study (felitti1998relationship).
+Depression is coded Y=1Y{=}1 if the answer to ’(Ever told) (you had) a depressive disorder (including depression, major depression, dysthymia, or minor depression)?’ is ’Yes’ and 0 otherwise.
+ACEs Item Questions Collapsed Encoding Method
+Any Sexual Abuse Number of times forced to touch, be touched, or have sex with anyone at least five years older or an adult Constructed by summing responses across multiple questions: all “never” encoded as 0, otherwise encoded as 1
+Household Substance Use Alcohol and illegal drug use in the household Constructed by summing responses across multiple questions: all “never” encoded as 0, otherwise encoded as 1
+Physical Abuse Frequency of physical abuse ‘Once’ or ‘More than once’ encoded as 1, ‘None’ encoded as 0
+Verbal Abuse Frequency of verbal abuse ‘Once’ or ‘More than once’ encoded as 1, ‘None’ encoded as 0
+Parental Divorce Response to parental divorce question ’Yes’ encoded as 1, ’No’ encoded as 0
+Living with a Depressed or Mentally Ill Person Response to question about living with a mentally ill person ’Yes’ encoded as 1, ’No’ encoded as 0
+Living with an Incarcerated Person Response to question about living with an incarcerated person ’Yes’ encoded as 1, ’No’ encoded as 0
+Physical Violence Between Parents Frequency of physical violence between parents ‘Once’ or ‘More than once’ encoded as 1, ‘None’ encoded as 0
+Emotional Neglect For how much of your childhood was there an adult in your household who made you feel safe and protected ‘Never’ and ‘A little of the time’ encoded as 1, ‘Some of the time’, ‘Most of the time’, ‘All of the time’ encoded as 0
+Physical Neglect For how much of your childhood was there an adult in your household who tried hard to make sure your basic needs were met ‘Never’ and ‘A little of the time’ encoded as 0, ‘Some of the time’, ‘Most of the time’, ‘All of the time’ encoded as 1
 Table 1. Data Processing Steps and Collapsed Indicators
@@
 Table 2. ACE Items in BRFSS 2023
-After processing the data, we restrict our sample to those who do not have any missing values in any ten ACEs variables.
+After processing the data, we restrict our sample to those who do not have any missing values in any of the ten ACE variables.
 We also omit those individuals who do not have the appropriate measurement for depression.
-For any missing family background, we directly use the imputation value provided by the BRFSS 2023.
-In total, there are 49,811 remaining individuals, among whom 23,506 individuals are from blue states [^2] Delaware, New Jersey, Oregon, Rhode Island, Virginia (as established in the 2024 United States presidential election) and 26,305 individuals from red states [^3] Florida, Georgia, Missouri, Nevada, Tennessee.
-In the part of data with individuals from blue states, 8,076 individuals do not have any ACEs while 15,512 were exposed to at least one ACE.
-In the part of data with individuals from blue states, 9,163 individuals do not have any ACEs while 17,231 were exposed to at least one ACE.
-3. Pre-specified Candidate Sets of ACEs
-We collect ten variables of ACEs from the BRFSS 2023 dataset.
-To examine the diverse effects of combinations of ACEs on depression, we first construct candidate sets of ACEs that we are interested in.
-Since previous literature has identified four ACEs as a critical threshold, beyond which the effects appear to intensify significantly, and that individuals with at least four ACEs are at an increased risk for chronic disease, poor mental health, and negative social outcomes (Hughes et al.
-[2017]; Waehrer et al.
-[2020]), we consider all quadruples of ACEs categories, all triples of ACEs categories, all pairs of ACEs categories, and all individual ACEs categories.
-We thus have 385 possible combinations of ACEs formed from ten candidate ACEs categories (see Appendix A).
-4. Stratification
-A previous study (Waehrer et al.
-[2020]) used BRFSS 2009 & 2012 to examine whether the relationship between ACEs and health outcomes, such as smoking, heavy drinking, and obesity, was similar across states, and controlled for baseline covariates including race/ethnicity, sex, and age.
-They also controlled for education, marital status and employment, which are post-treatment variables and likely to be affected by ACEs.
-Cox [1958] refers to these variables as “concomitant observations” and writes: “The essential point in our assumptions about these observations is that the value for any unit must be unaffected by the particular assignment of treatments to units actually used.
-In practice this means that either (a) the concomitant observations are taken before the assignment of treatments to units is made; or (b) the concomitant observations are made after the assignment of treatments, but before the effect of treatments has had time to develop…; or (c) we can assume from our knowledge of the nature of the concomitant observations concerned, that they are unaffected by the treatment.” If adjustments are made for variables that are affected by the treatment, then adjustments may remove part or all of the effect of the treatment, and hence estimators that adjust for a concomitant variable that has been affected by the treatment are generally biased (Rosenbaum [1984]).
-Therefore, we do not adjust for education, employment and marital status in our study.
-Given the relatively small number of covariates, we seek to mitigate potential confounding bias by stratifying participants on their recorded baseline characteristics (race/ethnicity, age, sex) and then comparing those exposed to specific ACE combinations (the “treated” group) with those reporting no ACE exposure (the “control” group) within each stratum.
-Each hypothesis contrasts a particular ACE combination against having no ACEs.
-Individuals who have any additional ACEs beyond the combination of interest remain in the treated group.
-For example, if our focus is on the effect of substance abuse, participants exposed to both substance abuse and physical abuse are still classified under “substance abuse.” We conduct these stratifications separately for individuals residing in blue states and those in red states, resulting in $385\times 2$ distinct stratified analyses.
-Specifically, we stratify participants by race/ethnicity, including White (non-Hispanic), Black (non-Hispanic), Hispanic, and Other (non-Hispanic); by age, with groups spanning 18-24, 25-34, 35-44, 45-54, 55-64, and 65+; and by sex, distinguishing between male and female participants.
-This results in a total of 48 distinct strata for both the blue and red subgroups.
-In the blue subgroup, the stratum with the lowest sample size contains 61 participants, with a median of 171.
-In the red subgroup, the stratum with the lowest sample size has 60 participants, with a median of 161.
-Detailed sample sizes for each stratum of both subgroups are provided in the appendix B.
+In total, there are 49,547 remaining individuals, among whom 23,390 individuals are from blue states [^2] Delaware, New Jersey, Oregon, Rhode Island, Virginia (as established in the 2024 United States presidential election) and 26,157 individuals from red states [^3] Florida, Georgia, Missouri, Nevada, Tennessee.
+In the part of data with individuals from blue states, 7,363 individuals do not have any ACEs while 16,027 were exposed to at least one ACE.
+In the part of data with individuals from red states, 8,479 individuals do not have any ACEs while 17,678 were exposed to at least one ACE.
+3. Isotonic Subgroup Selection
+We plan to use isotonic subgroup selection (ISS) to identify the high-risk group in our study (muller2025isotonic).
+Let $(X_{i},Y_{i})$, for $i=1,\dots,n$, be independent observations, where $X_{i}\in\mathbb{R}^{d}$ denotes covariates and $Y_{i}$ is the outcome of interest.
+In our setting, $Y_{i}$ is a binary indicator for depression, and $X_{i}$ consists of different combinations of ten ACE variables, each coded as binary.
+We assume a regression function $\eta(x)=\mathbb{E}(Y|X=x)$ and impose the coordinate-wise monotonicity assumption: $x\preceq x^{\prime}\implies\eta(x)\leq\eta(x^{\prime})$, where $x\preceq x^{\prime}$ means $x_{j}\leq x^{\prime}_{j}$ for all coordinates $j=1,\dots,d$.
+In practice, this monotonicity captures the expectation that exposure to additional ACEs should not reduce the probability of depression.
+For example, an individual reporting both physical abuse and parental substance use is assumed to have risk at least as high as someone reporting only parental substance use.
+Fix a threshold $\tau$.
+Our target is the superlevel region $X_{\tau}(\eta)=\{\,x:\eta(x)\geq\tau\,\}$, which we interpret as the subgroup of covariate profiles whose expected probability of depression exceeds $\tau$.
+The goal of Isotonic Subgroup Selection (ISS) is to estimate a subset $\hat{A}$ that approximates $X_{\tau}(\eta)$ while controlling error rates in a principled way.
+The ISS procedure consists of three main steps:
+  1. (1)
+    Hypothesis construction.
+    For each $X_{i}$, we test $H_{0,i}:\eta(X_{i})<\tau\quad\text{versus}\quad H_{1,i}:\eta(X_{i})\geq\tau.$ Because $Y$ is binary, classification-based $p$-values are constructed using anytime-valid martingale methods for bounded outcomes.
+Intuitively, these $p$-values are based on the cumulative evidence from neighboring profiles $X_{j}\preceq X_{i}$, comparing the observed number of depressed cases to what would be expected if $\eta(X_{i})<\tau$.
+This ensures that even with dependent tests across profiles, each $p_{i}$ is valid.
+    Formally, let $I(x)=\{\,j:X_{j}\preceq x\,\}$, $n(x)=|I(x)|$, and denote the concomitant responses by $Y_{(1)}(x),\dots,Y_{(n(x))}(x)$.
+Define $S_{k}(x)=\sum_{j=1}^{k}Y_{(j)}(x)$.
+Then an anytime-valid $p$-value for testing $\eta(x)<\tau$ is
+    $p_{\tau}(x)=\min_{1\leq k\leq n(x)}\frac{\tau^{\,S_{k}(x)}(1-\tau)^{\,k-S_{k}(x)+1}}{B(1-\tau;\,k-S_{k}(x)+1,\,S_{k}(x)+1)},$
+    where $B(z;a,b)$ is the incomplete Beta function. This construction ensures that $\mathbb{P}(p_{\tau}(x)\leq\alpha\mid X)\leq\alpha$ whenever $\eta(x)<\tau$ (see muller2025isotonic, Lemma 22).
+  2. (2)
+    Logical structure and multiple testing.
+    The set of hypotheses $\{H_{0,i}\}$ is not independent.
+By monotonicity, if $X_{i}\preceq X_{j}$, then $H_{0,j}\subseteq H_{0,i}$.
+To encode these logical relations, we construct a directed acyclic graph (DAG): $G=(I,E)$, where $I=\{1,\dots,m\}$ indexes the hypotheses.
+We include a directed edge $i\to j$ whenever $X_{j}\preceq X_{i}$, so that edges always point from a stronger profile to a weaker one.
+From $G$, we derive a polyforest $F$ by retaining at most one parent for each node.
+This choice is not unique and can be made, for example, at random, by selecting the closest parent in $\ell_{\infty}$ norm, or alternatively using other distance measures such as the $\ell_{2}$ norm.
+This sparsification reduces complexity while preserving logical consistency.
+    Let $R^{\text{ISS}}_{\alpha}(G,p)$ denote the rejection set from the DAG testing algorithm.
+The procedure begins by distributing the global $\alpha$-budget across the root nodes of $F$, with each root $i$ receiving share
+[…]
+    A root node $i$ is rejected if $p_{\tau,i}\leq\alpha_{i}$, in which case all its ancestors in $G$ are also rejected.
+It’s worth noting that here all its ancestors are defined with respect to the full DAG $G$, not only within the same tree of the polyforest.
+Any node that is logically stronger (i.e., dominates $i$ under the partial order) is simultaneously rejected.
+After each iteration, the rejected nodes are removed from $F$, and the unused $\alpha$ is reallocated to the remaining roots.
+The process continues until no further rejections are possible.
+(see formal pseudocode in Algorithm 1, muller2025isotonic)
+    Intuitively, the algorithm passes significance level downward through the hierarchy, where roots receive budget first, and rejected nodes release unused $\alpha$ to their descendants.
+Ancestors are rejected whenever a descendant is, ensuring monotonicity.
+The final rejection set $R^{\text{ISS}}_{\alpha}(G,p)$ therefore respects both the partial order and the global FWER.
+  3. (3)
+    Upward closure.
+    Denote the set of rejected nodes from the polyforest testing procedure as $R_{\alpha}^{\text{ISS}}(G,p)$.
+The estimated subgroup is then defined as $\hat{A}=\{\,x:x\succeq X_{i}\text{ for some }i\in R_{\alpha}^{\text{ISS}}(G,p)\,\}.$ By construction, $\hat{A}$ is the upward closure of $R_{\alpha}^{\text{ISS}}(G,p)$ with respect to the partial order $\preceq$.
+This closure step ensures that the selected subgroup respects the monotonicity structure: once a particular exposure profile $X_{i}$ is identified as high-risk, all more severe profiles $x\succeq X_{i}$ must also be included.
+Under this construction, ISS guarantees that with probability at least $1-\alpha$, no profile with $\eta(x)<\tau$ is included in $\hat{A}$, controlling the family-wise error rate (FWER).
+In applied terms, ISS provides a conservative but rigorous way to delineate ACE exposure patterns linked to elevated depression risk.
+Concretely, this means that the subgroups we identify consist only of ACE combinations whose estimated probability of depression exceeds the chosen threshold, while still capturing as many truly high-risk profiles as possible.
+In this way, ISS leverages both valid statistical guarantees and the monotonicity structure to yield interpretable subgroups of adverse childhood experiences most strongly associated with depression.
 4. Design of Data Analysis
-Our inference will be based on stratified sets of participants in the BRFSS 2023, as described previously in Section 4.
-We aim to explore the effect of different combinations of ACEs in our aforementioned candidate sets on depression and also identify those with higher risk.
-We will deem those combinations of ACEs for which the effect of the treatment is found to be significantly increase the odds of depression in both subgroups - those residing in blue states and red states - as replicable findings.
-Also, we are interested in those combinations more than doubling the odds ratio of depression in only one subgroup.
+We aim to explore the effect of different combinations of ACEs on depression and also identify those with higher risk.
+We will test all of the possible combinations of ACEs, therefore leading to 1024 ($2^{10}$) different hypotheses.
+We will deem those combinations of ACEs for which the effect more than doubles the odds of depression in both subgroups - those residing in blue states and red states - as replicable findings.
+Also, we are interested in those combinations more than doubling the odds ratio of depression in only one subgroup, that is testing the global null hypothesis.
 4.1. Data Turnover Framework
-We employ data turnover (Bekerman et al.
-in preparation, Bekerman et al.
-[2024a]) to analyze the data, ensuring control of the family-wise error rate (FWER; the chance of falsely rejecting any true null hypotheses at level $\alpha=0.05$) and facilitating flexible inference compared to automated cross-screening (Zhao et al.
-[2018]), by integrating both qualitative and quantitative insights from data exploration and formulating data-informed new hypotheses (e.g., Bekerman et al.
-[2024b]).
+We employ data turnover (bekerman2024protocol) to analyze the data, ensuring control of the family-wise error rate (FWER; the chance of falsely rejecting any true null hypotheses at level $\alpha=0.05$) and facilitating flexible inference compared to automated cross-screening (zhao2018cross), by integrating both qualitative and quantitative insights from data exploration and formulating data-informed new hypotheses (e.g., bekerman2024planning).
 A brief description is as follows.
@@
 In other words, both subgroups are used to screen and both subgroups are used to validate, but only one subgroup can generate novel hypotheses to be validated in the other subgroup while still maintaining the validity of our entire procedure.
-Because a limited number of combinations seek validation on the second subgroup, compared with directly controlling the FWER of 385 hypotheses on the whole data (eg.
-by using Bonferroni correction or Holm procedure), only a small fraction of the 385 potential hypotheses are considered in the validation sample.
+Because a limited number of combinations seek validation on the second subgroup, compared with directly controlling the FWER of 1024 hypotheses on the whole data (eg.
+by using Bonferroni correction, Holm procedure or directly applying ISS on the whole sample), only a small fraction of the 1024 potential hypotheses are considered in the validation sample.
+Moreover, we need to determine whether it is more beneficial to conduct EDA on the smaller or larger sample.
+Simulation results in Bekerman’s paper (bekerman2024protocol) suggest that performing EDA on the smaller subgroup—in our case, participants living in blue states—is advantageous.
+We anticipate that incorporating data exploration within this subgroup will strengthen our study’s overall design by enabling data-informed decision-making and facilitating the emergence of new hypotheses.
 4.2. Analysis Plan
-We are interested in testing whether a combination of ACEs would increase the odds of depression compared with the reference population.
-We plan to use the Cochran–Mantel–Haenszel test (CMH) in each subgroup to screen ACEs combinations.
-In the screening stage, we also need to define a cutoff of p-value (eg.
+As we note before, for each $X_{i}$, we test $H_{0,i}:\eta(X_{i})<\tau\quad\text{versus}\quad H_{1,i}:\eta(X_{i})\geq\tau.$ Therefore, we need to prespecify $\tau$.
+Intuitively, if $\tau$ is too small, the true superlevel set is very large and its corners tend to lie deep.
+If $\tau$ is too large, no method has power and all outputs are empty.
+So we are interested in testing whether a combination of ACEs would more than double the odds of depression compared with the reference population.
+Let $p_{0}$ be a baseline probability - proportion of people exposed to no ACEs.
+In previous BRFSS datasets which are not included in this study, $p_{0}$ is around $9.4\%$.
+To test that the odds at $x$ are at least $c$ times the baseline odds,
+$\frac{\tau}{1-\tau}=c\frac{p_{0}}{1-p_{0}}$
+$\tau\ =\frac{cp_{0}}{1+(c-1)p_{0}}$
+Use $c=2,p_{0}=9.4\%$, $\tau=0.172$.
+We plan to integrate isotonic subgroup selection (ISS) with the data turnover method.
+Recall that in data turnover framework, we split the data into two subgroups $\mathcal{I}_{R}$ (Red Subgroup) and $\mathcal{I}_{B}$ (Blue Subgroup).
+For every combination $X_{i}$ that we are interested, define the anytime-valid classification $p$-value $p_{\tau}(x)$ as in the binary setting (see Section 3).
+Below is the pre-specified analysis plan for the red subgroup.
+Compute red-side $p$-values $p^{R}_{\tau,i}:=p_{\tau}(X_{i};\mathcal{I}_{R})$ for all $i$, and form the screened index set $\mathcal{S}_{\kappa}=\bigl\{\,i:p^{R}_{\tau,i}\leq\kappa\,\bigr\}$ to be tested in $\mathcal{I}_{B}$, for a pre-specified screening cutoff $\kappa\in(0,1)$.
+From the full DAG $G$, we derive a polyforest $F$ by retaining at most one parent for each node, which can be made at random or based on a distance criterion.
+We could organize the polyforest $F_{\kappa}^{B}$ in $\mathcal{I}_{B}$ from the information ($p^{R}_{\tau,i}$, $\mathcal{S}_{\kappa}$) we obtain from the screening stage in $\mathcal{I}_{R}$.
+All nodes of the polyforest $F_{\kappa}^{B}$ are drawn from $\mathcal{S}_{\kappa}$.
+For each $i\in\mathcal{S}_{\kappa}$, define its set of covers
+$\mathrm{Cover}(i)\;=\;\bigl\{\,j\in\mathcal{S}_{\kappa}\,:\,X_{j}\succ X_{i}\ \text{and there is no }k\in\mathcal{S}_{\kappa}\text{ with }X_{j}\succ X_{k}\succ X_{i}\,\bigr\}.$
+Select at most one parent for $i$ from $\mathrm{Cover}(i)$ guided by $p$-value: $\mathrm{par}(i)\in\arg\min_{\,j\in\mathrm{Cover}(i)}\ p^{R}_{\tau,j}$, choosing among candidates with tied minimal p-value uniformly at random.
+If $\mathrm{Cover}(i)=\emptyset$, declare $i$ a root.
+This yields the polyforest $F_{\kappa}^{B}$ (a DAG with in-degree $\leq 1$).
+On $F_{\kappa}^{B}$, compute blue-side $p$-values only for screened nodes, $p^{B}_{\tau,i}:=p_{\tau}(X_{i};\mathcal{I}_{B})$ for $i\in\mathcal{S}_{\kappa}$, and run the ISS DAG-testing algorithm with polyforest $F_{\kappa}^{B}$ controlling the FWER for the blue part at level $\alpha/2$.
+Using $p_{\tau,i}^{R}$ to choose parents steers each node toward a path whose upstream evidence is strongest, so $\alpha$ is concentrated along promising chains.
+This mitigates cases where a large-$p$ maximal node would otherwise block the flow of $\alpha$ and prevent strong descendants from ever receiving budget, increasing the power.
+Figure 1 illustrates a case where the parent selection creates a bottleneck.
+Here, nodes $1$, 66, and $7$ are roots, each receiving part of the global $\alpha$-budget (shown in red).
+Although node $5$ has a small $p$-value ($0.01$), its designated parent is node $6$, whose $p$-value is relatively large ($0.1$).
+Since node $6$ is unlikely to be rejected with its assigned budget ($0.025$), the rejection cannot propagate to node $5$, and the signal at node $5$ is blocked.
+This demonstrates how attaching a promising child to an unpromising parent can prevent significant nodes from being discovered.
+Figure 1.
+DAG Testing: Example 1.
+Each node is labeled with its $p$-value (round brackets); in the induced polyforest-weighted DAG, solid arrows indicate edges that are retained in the polyforest (weight==1), while dashed arrows indicate edges that exist in the full DAG but are not retained in the polyforest (weight==0, since each node can have at most one parent); filled circles indicate hypotheses previously rejected.
+Panels (a)–(b) correspond to successive iterations of the algorithm, showing how rejections propagate through the structure.
+In contrast, Figure 2 shows the situation when node $5$ is attached through node $7$, whose $p$-value is smaller ($0.03$).
+After node $1$ is rejected in the first iteration ($p_{1}=0.01\leq 0.0125$), the unused budget is reallocated, giving node $7$ a larger share ($0.0333$).
+With $p_{7}=0.03$, node $7$ can then be rejected, releasing further budget to its child node $5$.
+Since node $5$ has $p_{5}=0.01$, it is subsequently rejected, and the rejection can continue to its descendants.
+Simultaneously, since node $6$ is the parent of node $5$ in the full DAG, even if its $p$-value is large, it can also be rejected in line with logical partial order.
+Compared to Figure 1, the $p$-guided parent selection in Figure 2 avoids the bottleneck at node $6$ and allows the procedure to detect the significant signal at node $5$.
+Figure 2.
+DAG Testing: Example 2.
+The setting is the same as in Figure 1, except that node $5$ is attached through node 77 rather than node $6$.
+Panels (a)–(d) correspond to successive iterations of the algorithm, showing how rejections propagate through the structure.
+In the screening stage, we also need to define the cutoff $\kappa$ of p-value (eg.
 0.05).
-In a paper defining multimorbidity (Silber et al.
-[2018]), they chose form a 99.9% confidence interval in checking if the interval for the odds ratio was completely above 2.
+In a paper defining multimorbidity (silber2018defining), they chose to form a 99.9% confidence interval in checking if the interval for the odds ratio was completely above 2.
 This is a stringent threshold to control for the number of hypotheses falsely passing the screening step.
 Yet, choosing this cutoff is rather heuristic given the sample sizes and number of hypotheses.
-A basic intuition would be: if the cutoff is too large, then almost all of the hypotheses would pass the screening stage, which would make this step meaningless; if the cutoff is too small, though only a few number of true nulls would falsely pass the screening stage, some true non-nulls would be screened out if the signals aren’t that strong.
-Thus, following the structure of the BRFSS 2023, we set the number of participants in each subgroup at 1000, and set the number of hypotheses at 385, and try several cutoffs of p-value (0.001, 0.01,0.025, 0.05, 0.1), also varying the number of replicable hypotheses (1,50,100,150,200,250,300,350,385) and the signal-to-noise ratio (a sequence from 0.5 to 5, with a common difference 0.1).
-The result (see Figure 1) shows that whatever the number of replicable hypotheses and signal-to-noise ratio, the power using 0.001 as the cutoff is always the lowest.
-But there’s only trivial difference of power between choosing 0.01,0.025, 0.05 or 0.1, and none of them consistently perform the best under different values of the number of replicable hypotheses and the signal-to-noise ratio.
-Figure 1. Simulation results of power with p-value cutoffs, number of replicable hypotheses and different signal-to-noise ratios.
-Note: the $x$ axis represents the signal-to-noise ratio, the the $y$ axis represents the power. The facet titles in the gray bars are numbers of replicable hypotheses.
+A basic intuition would be: if the cutoff is too large, then almost all of the hypotheses would pass the screening stage, which would make this step meaningless; if the cutoff is too small, though only a small number of true nulls would falsely pass the screening stage, some true non-nulls would be screened out if the signals aren’t that strong.
+In our testing procedure, when a root node is rejected, all its ancestors defined with respect to the full DAG are rejected.
+That is, even if the combination of exposures is logically stronger but yields a comparatively large $p$-value, while a weaker combination produces a small $p$-value and is rejected, the stronger combination will also be rejected.
+For example, suppose the subgroup defined by “parental substance use only” has $p=0.0001$ and is rejected at its $\alpha$ budget level, whereas the larger subgroup “parental substance use and physical abuse” has $p=0.40$.
+Because the second subgroup dominates the first in the partial order, it is rejected automatically once the first is, regardless of its own $p$-value.
+Thus, in our context, the cutoff $\kappa$ could be relatively more lenient.
 We intend to control the FWER at level $\alpha=0.05$.
-Therefore, we will apply Holm’s procedure (Holm [1979]) on each lists of p-values of selected hypotheses at level $\alpha/2=0.025$.
+Therefore, we will apply the DAG-testing procedure on each list of p-values of selected hypotheses at level $\alpha/2=0.025$.
 Accordingly, we choose 0.025 as our cutoff in the screening stage.
-In other words, if the p-value of one hypothesis given by the CMH test is smaller than 0.025, then this hypothesis will pass the screening step, otherwise, this will be screened out.
-After selecting a list of hypotheses from subgroup 1 from the candidate sets collection, we will test them on subgroup 2.
-Specifically, for each selected hypothesis, we calculate the p-value using the data from subgroup 2, in the same way as we did in the screening stage on subgroup 1, and utilize Holm’s method at 0.025 to validate.
-Then, we will create a list of hypotheses from subgroup 2, including those selected from the candidate sets collection and generated by EDA, and validate on subgroup 1.
-Moreover, we need to determine whether it is more beneficial to conduct EDA on the smaller or larger sample.
-Simulation results in Bekerman’s paper (in preparation) suggest that performing EDA on the smaller subgroup—in our case, participants living in blue states—is advantageous.
-We anticipate that incorporating data exploration within this subgroup will strengthen our study’s overall design by enabling data-informed decision-making and facilitating the emergence of new hypotheses.
+In other words, if the anytime-valid p-value of one hypothesis is smaller than 0.025, then this hypothesis will pass the screening step, otherwise, this will be screened out.
+For the blue subgroup - the one with smaller sample size, we will do exploratory data analysis (EDA).
+Of course, we could first do the pre-specified thing we plan on the red subgroup, but we have more flexibility on this part.
 This flexibility may allow us to identify more complex hypotheses or yield unexpected benefits.
-For example, we may explore alternative ways of grouping ACE categories, refine our criteria for encoding ACE items, or investigate potential effect modification.
-5.3. The high-risk group
-After employing data turnover, we define the group of combinations for which the global null is rejected, that is, at least in one subgroup, as the high-risk group; in other words, the relevant combinations of ACEs will be identified as placing individuals at higher risk of depression.
-The group is formed as the union of validated combinations from subgroup 1 and validated combinations from subgroup 2, and simultaneously, we can tell what combinations are replicable and are only validated on one subgroup.
-If any pre-specified hypotheses passes four tests - passes the screening stage in the first subgroup and the validation in the second subgroup, and also passes the screening stage in the second subgroup and the validation in the first subgroup - they will be deemed as replicable.
-For new hypotheses formulated in the EDA process, if they are validated in the other subgroup, related combinations will be in the high-risk group.
-We also provide a parsimonious list without duplicating any set whose components are already on the list as a smaller set.
-In other words, if a combination A in the collection is the subset of another combination B, we will remove B from the collection.
-For example, if both hypotheses regarding substance abuse and substance abuse + physical abuse are rejected, only substance abuse will be included in the parsimonious list.
-Previous literature has confirmed that ACEs can have a cumulative effect: each additional type of adversity elevates the probability of unfavorable outcomes (Felitti et al.
-[1998]; Chapman et al.
-[2004]).
-Therefore, given this monotonicity assumption, the validation of combination A implies that combination B increases the odds ratio of depression.
-In terms of replicable findings, other than combinations simultaneously rejected in both subgroups, we also include the combination of non-duplicate components from two different combinations respectively rejected in each subgroup, if this combination has not been contained, based on the monotonicity assumption.
-For example, if substance abuse is rejected on subgroup 1, but substance abuse + physical abuse is not (which does not make sense in real-world scenario but may happen due to some statistical reasons), and only substance abuse + physical abuse is rejected on subgroup 2, then substance abuse + physical abuse will be deemed replicable.
-The following tree diagrams (Figure 2, 3) makes things clear visually.
-We organize the ACE combinations in a tree structure, where each layer represents combinations of 1, 2, 3, or 4 ACE items, with each node corresponding to a unique combination.
-If a combination is rejected in the first subgroup, a yellow checkmark is placed at the top-left corner of the node, while a green checkmark appears at the bottom-right corner if the combination is rejected in the second subgroup.
-According to the monotonicity assumption, if a node is marked with either a yellow or green check, the entire subtree beneath that node inherits the same mark.
-Additionally, it’s important to note that identical combinations may appear in different subtrees within the same layer.
-If a combination node is marked in one subtree, all nodes with the same combination in other subtrees will be similarly marked.
-Ultimately, a combination node that is marked with checkmarks from both subgroups represents a replicable finding.
-In constructing the parsimonious list, we prioritize nodes from top to bottom, truncating any subtrees of replicable nodes as they are considered finalized, thus preventing redundant checks for combinations that have already been deemed replicable.
-In Example 1 (Figure 2), only substance abuse is rejected in subgroup 1, and this node is marked with a yellow checkmark.
-In contrast, only the combination substance abuse + physical abuse is rejected in subgroup 2, with a green checkmark placed on this node.
-Moving to the second layer, all nodes in the subtree of substance abuse inherit the yellow checkmark.
-As a result, the combination substance abuse + physical abuse is marked with both checkmarks and is considered replicable.
-Simultaneously, the subtrees of substance abuse + physical abuse are truncated, meaning that any combinations containing substance abuse and physical abuse are no longer considered for further evaluation.
-In Example 2 (Figure 3), only sexual abuse is rejected in subgroup 1, while only physical abuse is rejected in subgroup 2.
-Following the rules of placing checkmarks, sexual abuse + verbal abuse has two checkmarks and is deemed replicable.
-Also, the subtrees of sexual abuse + verbal abuse are truncated.
-Figure 2. Tree Diagram: Example 1
-Figure 3. Tree Diagram: Example 2
-Appendix A Candidate Sets of ACEs (variable names used)
-We consider all quadruples of ACEs categories, all triples of ACEs categories, all pairs of ACEs categories, and all individual ACEs category and form this candidate sets of ACEs.
-Table 3. Candidate sets of ACEs
-Item.1 Item.2 Item.3 Item.4
-1 ACESUB
-2 ACEPRISN
-3 ACEDIVRC
-4 ACEPUNCH
-5 ACEHURT1
-6 ACESWEAR
-7 ACESEX
-8 ACEADSAF
-9 ACEADNED
-10 ACEDEPRS
-11 ACESUB ACEPRISN
-12 ACESUB ACEDIVRC
-13 ACESUB ACEPUNCH
-14 ACESUB ACEHURT1
-15 ACESUB ACESWEAR
-16 ACESUB ACESEX
-17 ACESUB ACEADSAF
-18 ACESUB ACEADNED
-19 ACESUB ACEDEPRS
-20 ACEPRISN ACEDIVRC
-21 ACEPRISN ACEPUNCH
-22 ACEPRISN ACEHURT1
-23 ACEPRISN ACESWEAR
-24 ACEPRISN ACESEX
-25 ACEPRISN ACEADSAF
-26 ACEPRISN ACEADNED
-27 ACEPRISN ACEDEPRS
-28 ACEDIVRC ACEPUNCH
-29 ACEDIVRC ACEHURT1
-30 ACEDIVRC ACESWEAR
-31 ACEDIVRC ACESEX
-32 ACEDIVRC ACEADSAF
-33 ACEDIVRC ACEADNED
-34 ACEDIVRC ACEDEPRS
-35 ACEPUNCH ACEHURT1
-36 ACEPUNCH ACESWEAR
-37 ACEPUNCH ACESEX
-38 ACEPUNCH ACEADSAF
-39 ACEPUNCH ACEADNED
-40 ACEPUNCH ACEDEPRS
-41 ACEHURT1 ACESWEAR
-42 ACEHURT1 ACESEX
-43 ACEHURT1 ACEADSAF
-44 ACEHURT1 ACEADNED
-45 ACEHURT1 ACEDEPRS
-46 ACESWEAR ACESEX
-47 ACESWEAR ACEADSAF
-48 ACESWEAR ACEADNED
-49 ACESWEAR ACEDEPRS
-50 ACESEX ACEADSAF
-51 ACESEX ACEADNED
-52 ACESEX ACEDEPRS
-53 ACEADSAF ACEADNED
-54 ACEADSAF ACEDEPRS
-55 ACEADNED ACEDEPRS
-56 ACESUB ACEPRISN ACEDIVRC
-57 ACESUB ACEPRISN ACEPUNCH
-58 ACESUB ACEPRISN ACEHURT1
-59 ACESUB ACEPRISN ACESWEAR
-60 ACESUB ACEPRISN ACESEX
-61 ACESUB ACEPRISN ACEADSAF
-62 ACESUB ACEPRISN ACEADNED
-63 ACESUB ACEPRISN ACEDEPRS
-64 ACESUB ACEDIVRC ACEPUNCH
-65 ACESUB ACEDIVRC ACEHURT1
-66 ACESUB ACEDIVRC ACESWEAR
-67 ACESUB ACEDIVRC ACESEX
-68 ACESUB ACEDIVRC ACEADSAF
-69 ACESUB ACEDIVRC ACEADNED
-70 ACESUB ACEDIVRC ACEDEPRS
-71 ACESUB ACEPUNCH ACEHURT1
-72 ACESUB ACEPUNCH ACESWEAR
-73 ACESUB ACEPUNCH ACESEX
-74 ACESUB ACEPUNCH ACEADSAF
-75 ACESUB ACEPUNCH ACEADNED
-76 ACESUB ACEPUNCH ACEDEPRS
-77 ACESUB ACEHURT1 ACESWEAR
-78 ACESUB ACEHURT1 ACESEX
-79 ACESUB ACEHURT1 ACEADSAF
-80 ACESUB ACEHURT1 ACEADNED
-81 ACESUB ACEHURT1 ACEDEPRS
-82 ACESUB ACESWEAR ACESEX
-83 ACESUB ACESWEAR ACEADSAF
-84 ACESUB ACESWEAR ACEADNED
-85 ACESUB ACESWEAR ACEDEPRS
-86 ACESUB ACESEX ACEADSAF
-87 ACESUB ACESEX ACEADNED
-88 ACESUB ACESEX ACEDEPRS
-89 ACESUB ACEADSAF ACEADNED
-90 ACESUB ACEADSAF ACEDEPRS
-91 ACESUB ACEADNED ACEDEPRS
-92 ACEPRISN ACEDIVRC ACEPUNCH
-93 ACEPRISN ACEDIVRC ACEHURT1
-94 ACEPRISN ACEDIVRC ACESWEAR
-95 ACEPRISN ACEDIVRC ACESEX
-96 ACEPRISN ACEDIVRC ACEADSAF
-97 ACEPRISN ACEDIVRC ACEADNED
-98 ACEPRISN ACEDIVRC ACEDEPRS
-99 ACEPRISN ACEPUNCH ACEHURT1
-100 ACEPRISN ACEPUNCH ACESWEAR
-101 ACEPRISN ACEPUNCH ACESEX
-102 ACEPRISN ACEPUNCH ACEADSAF
-103 ACEPRISN ACEPUNCH ACEADNED
-104 ACEPRISN ACEPUNCH ACEDEPRS
-105 ACEPRISN ACEHURT1 ACESWEAR
-106 ACEPRISN ACEHURT1 ACESEX
-107 ACEPRISN ACEHURT1 ACEADSAF
-108 ACEPRISN ACEHURT1 ACEADNED
-109 ACEPRISN ACEHURT1 ACEDEPRS
-110 ACEPRISN ACESWEAR ACESEX
-111 ACEPRISN ACESWEAR ACEADSAF
-112 ACEPRISN ACESWEAR ACEADNED
-113 ACEPRISN ACESWEAR ACEDEPRS
-114 ACEPRISN ACESEX ACEADSAF
-115 ACEPRISN ACESEX ACEADNED
-116 ACEPRISN ACESEX ACEDEPRS
-117 ACEPRISN ACEADSAF ACEADNED
-118 ACEPRISN ACEADSAF ACEDEPRS
-119 ACEPRISN ACEADNED ACEDEPRS
-120 ACEDIVRC ACEPUNCH ACEHURT1
-121 ACEDIVRC ACEPUNCH ACESWEAR
-122 ACEDIVRC ACEPUNCH ACESEX
-123 ACEDIVRC ACEPUNCH ACEADSAF
-124 ACEDIVRC ACEPUNCH ACEADNED
-125 ACEDIVRC ACEPUNCH ACEDEPRS
-126 ACEDIVRC ACEHURT1 ACESWEAR
-127 ACEDIVRC ACEHURT1 ACESEX
-128 ACEDIVRC ACEHURT1 ACEADSAF
-129 ACEDIVRC ACEHURT1 ACEADNED
-130 ACEDIVRC ACEHURT1 ACEDEPRS
-131 ACEDIVRC ACESWEAR ACESEX
-132 ACEDIVRC ACESWEAR ACEADSAF
-133 ACEDIVRC ACESWEAR ACEADNED
-134 ACEDIVRC ACESWEAR ACEDEPRS
-135 ACEDIVRC ACESEX ACEADSAF
-136 ACEDIVRC ACESEX ACEADNED
-137 ACEDIVRC ACESEX ACEDEPRS
-138 ACEDIVRC ACEADSAF ACEADNED
-139 ACEDIVRC ACEADSAF ACEDEPRS
-140 ACEDIVRC ACEADNED ACEDEPRS
-141 ACEPUNCH ACEHURT1 ACESWEAR
-142 ACEPUNCH ACEHURT1 ACESEX
-143 ACEPUNCH ACEHURT1 ACEADSAF
-144 ACEPUNCH ACEHURT1 ACEADNED
-145 ACEPUNCH ACEHURT1 ACEDEPRS
-146 ACEPUNCH ACESWEAR ACESEX
-147 ACEPUNCH ACESWEAR ACEADSAF
-148 ACEPUNCH ACESWEAR ACEADNED
-149 ACEPUNCH ACESWEAR ACEDEPRS
-150 ACEPUNCH ACESEX ACEADSAF
-151 ACEPUNCH ACESEX ACEADNED
-152 ACEPUNCH ACESEX ACEDEPRS
-153 ACEPUNCH ACEADSAF ACEADNED
-154 ACEPUNCH ACEADSAF ACEDEPRS
-155 ACEPUNCH ACEADNED ACEDEPRS
-156 ACEHURT1 ACESWEAR ACESEX
-157 ACEHURT1 ACESWEAR ACEADSAF
-158 ACEHURT1 ACESWEAR ACEADNED
-159 ACEHURT1 ACESWEAR ACEDEPRS
-160 ACEHURT1 ACESEX ACEADSAF
-161 ACEHURT1 ACESEX ACEADNED
-162 ACEHURT1 ACESEX ACEDEPRS
-163 ACEHURT1 ACEADSAF ACEADNED
-164 ACEHURT1 ACEADSAF ACEDEPRS
-165 ACEHURT1 ACEADNED ACEDEPRS
-166 ACESWEAR ACESEX ACEADSAF
-167 ACESWEAR ACESEX ACEADNED
-168 ACESWEAR ACESEX ACEDEPRS
-169 ACESWEAR ACEADSAF ACEADNED
-170 ACESWEAR ACEADSAF ACEDEPRS
-171 ACESWEAR ACEADNED ACEDEPRS
-172 ACESEX ACEADSAF ACEADNED
-173 ACESEX ACEADSAF ACEDEPRS
-174 ACESEX ACEADNED ACEDEPRS
-175 ACEADSAF ACEADNED ACEDEPRS
-176 ACESUB ACEPRISN ACEDIVRC ACEPUNCH
-177 ACESUB ACEPRISN ACEDIVRC ACEHURT1
-178 ACESUB ACEPRISN ACEDIVRC ACESWEAR
-179 ACESUB ACEPRISN ACEDIVRC ACESEX
-180 ACESUB ACEPRISN ACEDIVRC ACEADSAF
-181 ACESUB ACEPRISN ACEDIVRC ACEADNED
-182 ACESUB ACEPRISN ACEDIVRC ACEDEPRS
-183 ACESUB ACEPRISN ACEPUNCH ACEHURT1
-184 ACESUB ACEPRISN ACEPUNCH ACESWEAR
-185 ACESUB ACEPRISN ACEPUNCH ACESEX
-186 ACESUB ACEPRISN ACEPUNCH ACEADSAF
-187 ACESUB ACEPRISN ACEPUNCH ACEADNED
-188 ACESUB ACEPRISN ACEPUNCH ACEDEPRS
-189 ACESUB ACEPRISN ACEHURT1 ACESWEAR
-190 ACESUB ACEPRISN ACEHURT1 ACESEX
-191 ACESUB ACEPRISN ACEHURT1 ACEADSAF
-192 ACESUB ACEPRISN ACEHURT1 ACEADNED
-193 ACESUB ACEPRISN ACEHURT1 ACEDEPRS
-194 ACESUB ACEPRISN ACESWEAR ACESEX
-195 ACESUB ACEPRISN ACESWEAR ACEADSAF
-196 ACESUB ACEPRISN ACESWEAR ACEADNED
-197 ACESUB ACEPRISN ACESWEAR ACEDEPRS
-198 ACESUB ACEPRISN ACESEX ACEADSAF
-199 ACESUB ACEPRISN ACESEX ACEADNED
-200 ACESUB ACEPRISN ACESEX ACEDEPRS
-201 ACESUB ACEPRISN ACEADSAF ACEADNED
-202 ACESUB ACEPRISN ACEADSAF ACEDEPRS
-203 ACESUB ACEPRISN ACEADNED ACEDEPRS
-204 ACESUB ACEDIVRC ACEPUNCH ACEHURT1
-205 ACESUB ACEDIVRC ACEPUNCH ACESWEAR
-206 ACESUB ACEDIVRC ACEPUNCH ACESEX
-207 ACESUB ACEDIVRC ACEPUNCH ACEADSAF
-208 ACESUB ACEDIVRC ACEPUNCH ACEADNED
-209 ACESUB ACEDIVRC ACEPUNCH ACEDEPRS
-210 ACESUB ACEDIVRC ACEHURT1 ACESWEAR
-211 ACESUB ACEDIVRC ACEHURT1 ACESEX
-212 ACESUB ACEDIVRC ACEHURT1 ACEADSAF
-213 ACESUB ACEDIVRC ACEHURT1 ACEADNED
-214 ACESUB ACEDIVRC ACEHURT1 ACEDEPRS
-215 ACESUB ACEDIVRC ACESWEAR ACESEX
-216 ACESUB ACEDIVRC ACESWEAR ACEADSAF
-217 ACESUB ACEDIVRC ACESWEAR ACEADNED
-218 ACESUB ACEDIVRC ACESWEAR ACEDEPRS
-219 ACESUB ACEDIVRC ACESEX ACEADSAF
-220 ACESUB ACEDIVRC ACESEX ACEADNED
-221 ACESUB ACEDIVRC ACESEX ACEDEPRS
-222 ACESUB ACEDIVRC ACEADSAF ACEADNED
-223 ACESUB ACEDIVRC ACEADSAF ACEDEPRS
-224 ACESUB ACEDIVRC ACEADNED ACEDEPRS
-225 ACESUB ACEPUNCH ACEHURT1 ACESWEAR
-226 ACESUB ACEPUNCH ACEHURT1 ACESEX
-227 ACESUB ACEPUNCH ACEHURT1 ACEADSAF
-228 ACESUB ACEPUNCH ACEHURT1 ACEADNED
-229 ACESUB ACEPUNCH ACEHURT1 ACEDEPRS
-230 ACESUB ACEPUNCH ACESWEAR ACESEX
-231 ACESUB ACEPUNCH ACESWEAR ACEADSAF
-232 ACESUB ACEPUNCH ACESWEAR ACEADNED
-233 ACESUB ACEPUNCH ACESWEAR ACEDEPRS
-234 ACESUB ACEPUNCH ACESEX ACEADSAF
-235 ACESUB ACEPUNCH ACESEX ACEADNED
-236 ACESUB ACEPUNCH ACESEX ACEDEPRS
-237 ACESUB ACEPUNCH ACEADSAF ACEADNED
-238 ACESUB ACEPUNCH ACEADSAF ACEDEPRS
-239 ACESUB ACEPUNCH ACEADNED ACEDEPRS
-240 ACESUB ACEHURT1 ACESWEAR ACESEX
-241 ACESUB ACEHURT1 ACESWEAR ACEADSAF
-242 ACESUB ACEHURT1 ACESWEAR ACEADNED
-243 ACESUB ACEHURT1 ACESWEAR ACEDEPRS
-244 ACESUB ACEHURT1 ACESEX ACEADSAF
-245 ACESUB ACEHURT1 ACESEX ACEADNED
-246 ACESUB ACEHURT1 ACESEX ACEDEPRS
-247 ACESUB ACEHURT1 ACEADSAF ACEADNED
-248 ACESUB ACEHURT1 ACEADSAF ACEDEPRS
-249 ACESUB ACEHURT1 ACEADNED ACEDEPRS
-250 ACESUB ACESWEAR ACESEX ACEADSAF
-251 ACESUB ACESWEAR ACESEX ACEADNED
-252 ACESUB ACESWEAR ACESEX ACEDEPRS
-253 ACESUB ACESWEAR ACEADSAF ACEADNED
-254 ACESUB ACESWEAR ACEADSAF ACEDEPRS
-255 ACESUB ACESWEAR ACEADNED ACEDEPRS
-256 ACESUB ACESEX ACEADSAF ACEADNED
-257 ACESUB ACESEX ACEADSAF ACEDEPRS
-258 ACESUB ACESEX ACEADNED ACEDEPRS
-259 ACESUB ACEADSAF ACEADNED ACEDEPRS
-260 ACEPRISN ACEDIVRC ACEPUNCH ACEHURT1
-261 ACEPRISN ACEDIVRC ACEPUNCH ACESWEAR
-262 ACEPRISN ACEDIVRC ACEPUNCH ACESEX
-263 ACEPRISN ACEDIVRC ACEPUNCH ACEADSAF
-264 ACEPRISN ACEDIVRC ACEPUNCH ACEADNED
-265 ACEPRISN ACEDIVRC ACEPUNCH ACEDEPRS
-266 ACEPRISN ACEDIVRC ACEHURT1 ACESWEAR
-267 ACEPRISN ACEDIVRC ACEHURT1 ACESEX
-268 ACEPRISN ACEDIVRC ACEHURT1 ACEADSAF
-269 ACEPRISN ACEDIVRC ACEHURT1 ACEADNED
-270 ACEPRISN ACEDIVRC ACEHURT1 ACEDEPRS
-271 ACEPRISN ACEDIVRC ACESWEAR ACESEX
-272 ACEPRISN ACEDIVRC ACESWEAR ACEADSAF
-273 ACEPRISN ACEDIVRC ACESWEAR ACEADNED
-274 ACEPRISN ACEDIVRC ACESWEAR ACEDEPRS
-275 ACEPRISN ACEDIVRC ACESEX ACEADSAF
-276 ACEPRISN ACEDIVRC ACESEX ACEADNED
-277 ACEPRISN ACEDIVRC ACESEX ACEDEPRS
-278 ACEPRISN ACEDIVRC ACEADSAF ACEADNED
-279 ACEPRISN ACEDIVRC ACEADSAF ACEDEPRS
-280 ACEPRISN ACEDIVRC ACEADNED ACEDEPRS
-281 ACEPRISN ACEPUNCH ACEHURT1 ACESWEAR
-282 ACEPRISN ACEPUNCH ACEHURT1 ACESEX
-283 ACEPRISN ACEPUNCH ACEHURT1 ACEADSAF
-284 ACEPRISN ACEPUNCH ACEHURT1 ACEADNED
-285 ACEPRISN ACEPUNCH ACEHURT1 ACEDEPRS
-286 ACEPRISN ACEPUNCH ACESWEAR ACESEX
-287 ACEPRISN ACEPUNCH ACESWEAR ACEADSAF
-288 ACEPRISN ACEPUNCH ACESWEAR ACEADNED
-289 ACEPRISN ACEPUNCH ACESWEAR ACEDEPRS
-290 ACEPRISN ACEPUNCH ACESEX ACEADSAF
-291 ACEPRISN ACEPUNCH ACESEX ACEADNED
-292 ACEPRISN ACEPUNCH ACESEX ACEDEPRS
-293 ACEPRISN ACEPUNCH ACEADSAF ACEADNED
-294 ACEPRISN ACEPUNCH ACEADSAF ACEDEPRS
-295 ACEPRISN ACEPUNCH ACEADNED ACEDEPRS
-296 ACEPRISN ACEHURT1 ACESWEAR ACESEX
-297 ACEPRISN ACEHURT1 ACESWEAR ACEADSAF
-298 ACEPRISN ACEHURT1 ACESWEAR ACEADNED
-299 ACEPRISN ACEHURT1 ACESWEAR ACEDEPRS
-300 ACEPRISN ACEHURT1 ACESEX ACEADSAF
-301 ACEPRISN ACEHURT1 ACESEX ACEADNED
-302 ACEPRISN ACEHURT1 ACESEX ACEDEPRS
-303 ACEPRISN ACEHURT1 ACEADSAF ACEADNED
-304 ACEPRISN ACEHURT1 ACEADSAF ACEDEPRS
-305 ACEPRISN ACEHURT1 ACEADNED ACEDEPRS
-306 ACEPRISN ACESWEAR ACESEX ACEADSAF
-307 ACEPRISN ACESWEAR ACESEX ACEADNED
-308 ACEPRISN ACESWEAR ACESEX ACEDEPRS
-309 ACEPRISN ACESWEAR ACEADSAF ACEADNED
-310 ACEPRISN ACESWEAR ACEADSAF ACEDEPRS
-311 ACEPRISN ACESWEAR ACEADNED ACEDEPRS
-312 ACEPRISN ACESEX ACEADSAF ACEADNED
-313 ACEPRISN ACESEX ACEADSAF ACEDEPRS
-314 ACEPRISN ACESEX ACEADNED ACEDEPRS
-315 ACEPRISN ACEADSAF ACEADNED ACEDEPRS
-316 ACEDIVRC ACEPUNCH ACEHURT1 ACESWEAR
-317 ACEDIVRC ACEPUNCH ACEHURT1 ACESEX
-318 ACEDIVRC ACEPUNCH ACEHURT1 ACEADSAF
-319 ACEDIVRC ACEPUNCH ACEHURT1 ACEADNED
-320 ACEDIVRC ACEPUNCH ACEHURT1 ACEDEPRS
-321 ACEDIVRC ACEPUNCH ACESWEAR ACESEX
-322 ACEDIVRC ACEPUNCH ACESWEAR ACEADSAF
-323 ACEDIVRC ACEPUNCH ACESWEAR ACEADNED
-324 ACEDIVRC ACEPUNCH ACESWEAR ACEDEPRS
-325 ACEDIVRC ACEPUNCH ACESEX ACEADSAF
-326 ACEDIVRC ACEPUNCH ACESEX ACEADNED
-327 ACEDIVRC ACEPUNCH ACESEX ACEDEPRS
-328 ACEDIVRC ACEPUNCH ACEADSAF ACEADNED
-329 ACEDIVRC ACEPUNCH ACEADSAF ACEDEPRS
-330 ACEDIVRC ACEPUNCH ACEADNED ACEDEPRS
-331 ACEDIVRC ACEHURT1 ACESWEAR ACESEX
-332 ACEDIVRC ACEHURT1 ACESWEAR ACEADSAF
-333 ACEDIVRC ACEHURT1 ACESWEAR ACEADNED
-334 ACEDIVRC ACEHURT1 ACESWEAR ACEDEPRS
-335 ACEDIVRC ACEHURT1 ACESEX ACEADSAF
-336 ACEDIVRC ACEHURT1 ACESEX ACEADNED
-337 ACEDIVRC ACEHURT1 ACESEX ACEDEPRS
-338 ACEDIVRC ACEHURT1 ACEADSAF ACEADNED
-339 ACEDIVRC ACEHURT1 ACEADSAF ACEDEPRS
-340 ACEDIVRC ACEHURT1 ACEADNED ACEDEPRS
-341 ACEDIVRC ACESWEAR ACESEX ACEADSAF
-342 ACEDIVRC ACESWEAR ACESEX ACEADNED
-343 ACEDIVRC ACESWEAR ACESEX ACEDEPRS
-344 ACEDIVRC ACESWEAR ACEADSAF ACEADNED
-345 ACEDIVRC ACESWEAR ACEADSAF ACEDEPRS
-346 ACEDIVRC ACESWEAR ACEADNED ACEDEPRS
-347 ACEDIVRC ACESEX ACEADSAF ACEADNED
-348 ACEDIVRC ACESEX ACEADSAF ACEDEPRS
-349 ACEDIVRC ACESEX ACEADNED ACEDEPRS
-350 ACEDIVRC ACEADSAF ACEADNED ACEDEPRS
-351 ACEPUNCH ACEHURT1 ACESWEAR ACESEX
-352 ACEPUNCH ACEHURT1 ACESWEAR ACEADSAF
-353 ACEPUNCH ACEHURT1 ACESWEAR ACEADNED
-354 ACEPUNCH ACEHURT1 ACESWEAR ACEDEPRS
-355 ACEPUNCH ACEHURT1 ACESEX ACEADSAF
-356 ACEPUNCH ACEHURT1 ACESEX ACEADNED
-357 ACEPUNCH ACEHURT1 ACESEX ACEDEPRS
-358 ACEPUNCH ACEHURT1 ACEADSAF ACEADNED
-359 ACEPUNCH ACEHURT1 ACEADSAF ACEDEPRS
-360 ACEPUNCH ACEHURT1 ACEADNED ACEDEPRS
-361 ACEPUNCH ACESWEAR ACESEX ACEADSAF
-362 ACEPUNCH ACESWEAR ACESEX ACEADNED
-363 ACEPUNCH ACESWEAR ACESEX ACEDEPRS
-364 ACEPUNCH ACESWEAR ACEADSAF ACEADNED
-365 ACEPUNCH ACESWEAR ACEADSAF ACEDEPRS
-366 ACEPUNCH ACESWEAR ACEADNED ACEDEPRS
-367 ACEPUNCH ACESEX ACEADSAF ACEADNED
-368 ACEPUNCH ACESEX ACEADSAF ACEDEPRS
-369 ACEPUNCH ACESEX ACEADNED ACEDEPRS
-370 ACEPUNCH ACEADSAF ACEADNED ACEDEPRS
-371 ACEHURT1 ACESWEAR ACESEX ACEADSAF
-372 ACEHURT1 ACESWEAR ACESEX ACEADNED
-373 ACEHURT1 ACESWEAR ACESEX ACEDEPRS
-374 ACEHURT1 ACESWEAR ACEADSAF ACEADNED
-375 ACEHURT1 ACESWEAR ACEADSAF ACEDEPRS
-376 ACEHURT1 ACESWEAR ACEADNED ACEDEPRS
-377 ACEHURT1 ACESEX ACEADSAF ACEADNED
-378 ACEHURT1 ACESEX ACEADSAF ACEDEPRS
-379 ACEHURT1 ACESEX ACEADNED ACEDEPRS
-380 ACEHURT1 ACEADSAF ACEADNED ACEDEPRS
-381 ACESWEAR ACESEX ACEADSAF ACEADNED
-382 ACESWEAR ACESEX ACEADSAF ACEDEPRS
-383 ACESWEAR ACESEX ACEADNED ACEDEPRS
-384 ACESWEAR ACEADSAF ACEADNED ACEDEPRS
-385 ACESEX ACEADSAF ACEADNED ACEDEPRS
-Appendix B Sample Sizes by Stratum in Each Subgroup
-Table 4. Sample Sizes by Stratum in the Blue Subgroup
-Race/Ethnicity Sex Age Sample Size
-1 White, Non-Hispanic Male 18-24 362
-2 Black, Non-Hispanic Male 18-24 61
-3 Hispanic Male 18-24 182
-4 Other race, Non-Hispanic Male 18-24 91
-5 White, Non-Hispanic Female 18-24 232
-6 Black, Non-Hispanic Female 18-24 63
-7 Hispanic Female 18-24 135
-8 Other race, Non-Hispanic Female 18-24 72
-9 White, Non-Hispanic Male 25-34 695
-10 Black, Non-Hispanic Male 25-34 82
-11 Hispanic Male 25-34 234
-12 Other race, Non-Hispanic Male 25-34 133
-13 White, Non-Hispanic Female 25-34 621
-14 Black, Non-Hispanic Female 25-34 99
-15 Hispanic Female 25-34 274
-16 Other race, Non-Hispanic Female 25-34 125
-17 White, Non-Hispanic Male 35-44 820
-18 Black, Non-Hispanic Male 35-44 103
-19 Hispanic Male 35-44 222
-20 Other race, Non-Hispanic Male 35-44 194
-21 White, Non-Hispanic Female 35-44 908
-22 Black, Non-Hispanic Female 35-44 157
-23 Hispanic Female 35-44 247
-24 Other race, Non-Hispanic Female 35-44 129
-25 White, Non-Hispanic Male 45-54 955
-26 Black, Non-Hispanic Male 45-54 135
-27 Hispanic Male 45-54 171
-28 Other race, Non-Hispanic Male 45-54 131
-29 White, Non-Hispanic Female 45-54 1212
-30 Black, Non-Hispanic Female 45-54 168
-31 Hispanic Female 45-54 206
-32 Other race, Non-Hispanic Female 45-54 151
-33 White, Non-Hispanic Male 55-64 1513
-34 Black, Non-Hispanic Male 55-64 141
-35 Hispanic Male 55-64 160
-36 Other race, Non-Hispanic Male 55-64 132
-37 White, Non-Hispanic Female 55-64 2054
-38 Black, Non-Hispanic Female 55-64 267
-39 Hispanic Female 55-64 171
-40 Other race, Non-Hispanic Female 55-64 110
-41 White, Non-Hispanic Male 65+ 3688
-42 Black, Non-Hispanic Male 65+ 260
-43 Hispanic Male 65+ 118
-44 Other race, Non-Hispanic Male 65+ 180
-45 White, Non-Hispanic Female 65+ 4561
-46 Black, Non-Hispanic Female 65+ 473
-47 Hispanic Female 65+ 139
-48 Other race, Non-Hispanic Female 65+ 169
-Table 5. Sample Sizes by Stratum in the Red Subgroup
-Race/Ethnicity Sex Age Sample Size
-1 White, Non-Hispanic Male 18-24 362
-2 Black, Non-Hispanic Male 18-24 61
-3 Hispanic Male 18-24 182
-4 Other race, Non-Hispanic Male 18-24 91
-5 White, Non-Hispanic Female 18-24 232
-6 Black, Non-Hispanic Female 18-24 63
-7 Hispanic Female 18-24 135
-8 Other race, Non-Hispanic Female 18-24 72
-9 White, Non-Hispanic Male 25-34 695
-10 Black, Non-Hispanic Male 25-34 82
-11 Hispanic Male 25-34 234
-12 Other race, Non-Hispanic Male 25-34 133
-13 White, Non-Hispanic Female 25-34 621
-14 Black, Non-Hispanic Female 25-34 99
-15 Hispanic Female 25-34 274
-16 Other race, Non-Hispanic Female 25-34 125
-17 White, Non-Hispanic Male 35-44 820
-18 Black, Non-Hispanic Male 35-44 103
-19 Hispanic Male 35-44 222
-20 Other race, Non-Hispanic Male 35-44 194
-21 White, Non-Hispanic Female 35-44 908
-22 Black, Non-Hispanic Female 35-44 157
-23 Hispanic Female 35-44 247
-24 Other race, Non-Hispanic Female 35-44 129
-25 White, Non-Hispanic Male 45-54 955
-26 Black, Non-Hispanic Male 45-54 135
-27 Hispanic Male 45-54 171
-28 Other race, Non-Hispanic Male 45-54 131
-29 White, Non-Hispanic Female 45-54 1212
-30 Black, Non-Hispanic Female 45-54 168
-31 Hispanic Female 45-54 206
-32 Other race, Non-Hispanic Female 45-54 151
-33 White, Non-Hispanic Male 55-64 1513
-34 Black, Non-Hispanic Male 55-64 141
-35 Hispanic Male 55-64 160
-36 Other race, Non-Hispanic Male 55-64 132
-37 White, Non-Hispanic Female 55-64 2054
-38 Black, Non-Hispanic Female 55-64 267
-39 Hispanic Female 55-64 171
-40 Other race, Non-Hispanic Female 55-64 110
-41 White, Non-Hispanic Male 65+ 3688
-42 Black, Non-Hispanic Male 65+ 260
-43 Hispanic Male 65+ 118
-44 Other race, Non-Hispanic Male 65+ 180
-45 White, Non-Hispanic Female 65+ 4561
-46 Black, Non-Hispanic Female 65+ 473
-47 Hispanic Female 65+ 139
-48 Other race, Non-Hispanic Female 65+ 169
+For example, we may explore alternative ways of grouping ACE categories and refine our criteria for encoding ACE items.